A Selected Bibliography of Publications by, and about, Otto Robert Frisch

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

28 March 2025 Version 1.17

Title word cross-reference

$29.95 [Sta08]. 2^{3}P_{0,1,2} [FP31]. 2^{3}S_{1} [FP31]. $30 [Sta08]. $32 [Sta08]. $7.50 [Lin60]. ^{235} [SFFdH45]. α [Fri42b].

-ray [Fri42b].

1 [FP64, SBRT^{+}16a]. 114 [Fri61d]. 11th [AP77]. 12th [AP77]. 1904-1979 [ALJ82]. 1930s [Hen80, Stu79]. 1939 [MF69]. 1940 [Arn03]. 1954 [Rot54]. 1960s [Mla98]. 1963 [Gar63a]. 1963-A [Gar63a]. 1968 [Fri70b]. 1979 [Fel79, Pei81]. 1995 [Lee07]. 1999 [Rot00].

2 [SBRT^{+}16b]. 2002 [Bar05]. 2010 [KLR13]. 20th [FPLR59a, Fri73b, Rei07]. 272 [Fri73a]. 285 [Lip59].

3 [For17, SBRT^{+}16c]. 310 [But64]. 368 [Fri73b].

4 [SBRT^{+}16d]. 448pp [For17].

737 [Fri74a].

80th [FPLR59b, Lin60]. 8th [Rot00].

93 [Nod34]. 978-0-19-874430-6 [For17].

A.E.R.E. [Fri47a]. Ablenkung [FS33e]. Abraham [Smo07]. Absolute [Fri42a]. Absorption [FHM36, FP43]. Academy [Bar05]. Acceleration [Fri51a, Fri51c, Fri50a]. accelerators [Fri53b, Fri55f]. act [Stu13]. acting [FvK37b]. adventures [FL06]. af [BR64, RRK^{+}64]. after [Coc46, Kae48, SFFdH45]. Age [Stu18, Fri59l, Fri79b, Rif06]. ago [Sim14]. Alamos [McC24, SR92, Sza92a, Sza92b]. Albert [Smo07, Fri79c, Fri73a]. alkaline [HS39]. Allen [Fri55d]. always [Seg93]. America [Stu85]. Amherst [Bar05]. amplifiers [FF46]. Amrine [Fri60b]. analyser [FFS51, FO60]. Analysis [GJ23, KED^{+}60, Fri42b]. Analysis/Mehlin [KED^{+}60]. Anderson [Fri62b]. angular [Fri30b]. Announced [Ano39]. Anomalien [Fri33a, FS33d]. Anomalies [Fri33a, FS33d]. Anti [Fri56b, Fri55c, Fri56c, Fri56d, Fri57g]. anti-matter [Fri56b, Fri57g]. Anti-particles [Fri56b]. anti-proton [Fri55c, Fri56c, Fri56d]. anything [Fri79f]. Apology [Pra93]. application [Fri59e]. Archives [Bar05]. areas [FHS31]. Arising [Bas71, Bas09]. Arms [BGR61, BGR61]. Arthur [BGR61]. Artificial [Fri51a, Sim07, Fri51c, Fri55a]. artificielle [Fri55a]. Astronomy [KED^{+}60, Gar63a]. Astronomy/Born [KED^{+}60]. Atom [Ano39, Fri68b, Kae45, Lau51, Smi61, Gol87, Bro47, GHJ47]. atome [Gol87]. Atomic [BBD^{+}46, Boh61, Bro47, FR13, Fri51a, Fri51b, Fri54h, Fri54b, Fri54i, Fri54c, Fri55d, Fri61c, GHJ47, Gol90, Gow64, Kel04, Lau51, Lin60, Pei97, Rot54, WM15, WHF^{+}96, Arn03, FP39, Fri41, Fri43, Fri51c, Fri54d, Fri55e, Fri57f, Fri59k, FPLR59b, Fri60c, Fri64g, Fri67b, Fri77c, Fri79b, GA74a, GA74b, OPM38, Ree15a, Ros98, SR92, Hal61, Smi61]. atomkernen [Fri64g]. Atoms [Bro47, Fri47b, Fri47c, Fri59g, Fri65d, Bro47, Fri51d, GHJ47, Fri55d]. Ausmessen [Fri28]. Autobiographical [Pra93]. Autobiographies [For17, GT15]. autobiography [Seg93]. automatic [DFS70, FO60]. aux [Per39]. average [FP31].

B [Fri73a]. background [Ree15b]. balance [Fri30b]. Band [SBRT^{+}16a, SBRT^{+}16b, SBRT^{+}16c, SBRT^{+}16d]. based [FP40c, FP64, FP65]. Battle [Kae45]. be [Fri57d]. Beams [Fri65a, EFS31, EFS32, FS32, Fri33a, FS33d, Fri59a]. became [Ree15a]. Began [FW67, Kae48, Fri54h]. Behind [Bar05]. being [Rot54]. Beiträge [KED^{+}60, FPLR59a, Lip59]. beliefs [DWS79]. Bellman [KED^{+}60]. Berlin [KLR13]. Bernard [Fri74a]. berücksichtigt [CSW97]. Bestrahlung [HS39]. Betrachtungen [Fri59c]. Between [Stu18]. Beugung [Fri33a, FS33d, FS31].

Beyond [Bas71, Bas09]. Big [Ano39]. Biographical [Bar05]. Biographie [Nef05, Nef06]. Biography [Nef05, Nef06, Nef07, Smo07, Sta08]. Biologists [Dei96]. birth [Fri61e]. birthday [FPLR59b, Lin60]. black [FP43]. Bohr [Fri67a, BR64, Fre86, Fri67a, RRK^{+}64, Roz67, Sch00]. Bomb [Fri68b, Gre00, Kae45, Kel04, Pea24, WM15, WHF^{+}96, Far13, FP40c, FP40b, FP40d, Fri54e, Fri54g, FP64, FP65, FP04, Ree15a, Ree24a, Ros98, SR92, Fri61e]. Bombardment [Fri39c, MF39b]. bombs [Arn03, McC25]. Book [But64, Fes54, Fir64, For17, Fri50a, Fri51d, Fri55d, Fri57b, Fri59e, Fri59d, Fri60b, Fri61e, Fri61d, Fri62b, Fri62c, Fri64a, Fri64e, Fri64d, Fri69, Fri73a, Fri73b, Fri74a, Fri77a, Hal61, Hen80, Lei68, Lin60, Lip59, Mit86, Pei79, Sei80, Sie80, Sta08, Wil80, Fri57a]. Books [Ano64, Bar05, BGR61, Bro47, Fri67a]. Born [KED^{+}60]. boundaries [Rei07]. Bragg [GT15, For17]. Braunschweig [Lip59]. breakthroughs [Lig05]. Bremsstrahlung [Fri59c, FO59]. bridging [Rei07]. Bringing [Stu85]. Britain [Arn00, Gow64, GA74a, GA74b]. British [Sza92a, Sza92b]. Broadcasting [Fri57b]. Broadway [Sch00]. Broglie [EFS31, EFS32, EFS31, EFS32]. Broglie-Wellen [EFS31, EFS32]. Brum [Gre00]. bubble [FO60, Fri72a, Fri73c]. bubble-chamber [Fri72a]. Bücher [KED^{+}60]. build [SR92].

C [BGR61, Fri59e, FvK37f]. C-Neutrons [FvK37f]. C.B.E. [AP77]. Calcium [Fei35]. Calcul [Per39]. Calculation [Fri39a, McC24, Per39]. calibration [Fri42a]. Cambridge [ALJ82, Bar05, Fri61d, Pei79, Coc46]. came [Fri57d]. Campbell [Bro47, GHJ47]. can [Fri79f]. Cancer [Kae39]. Cannonball [Ano39]. Capek [Fri62c]. Capture [FP36, FvK37a, FvK37e, Fri37]. Carter [Smo07]. cathode [Fri26, Fri27]. Causality [Fri53a, Fri64f]. Census [Bar05]. Century [Bar05, FPLR59a, Lig05, Rei07, Fri73b]. Cerenkov [Fri59e]. chaîne [Per39]. Chain [FP40c, FP64, FP65, And73, Fri77b, Per39]. chamber [FO60, Fri72a, Fri73c]. Change [Mit86]. changed [Ber11]. changer [Ree15a]. character [FL06]. Characteristics [And98]. Chemical [Rei07]. Chemie [FPLR59a, KED^{+}60, Lip59]. Chemistry [Lei68, FPLR59a, Gar63a, Bar05]. Churchill [Far13]. Churchman [KED^{+}60]. circle [BR64]. Clark [Fri61e]. Classic [FL06]. Classics [Pra93]. Claudio [Mit86]. Clear [For17, GT15]. Clocks [Smo07]. cloth [Sta08]. co [RRK^{+}64]. co-workers [RRK^{+}64]. Coal [Fri54a, Fri56f]. Coal-Driven [Fri54a, Fri56f]. Coherent [FO59, Fri59c]. colleagues [BR64, Roz67]. Collected [Fri74a, Smo07]. collects [Fri64g]. College [ALJ82]. Colloquium [Bas71, Bas09]. Columbia [Ano39, Lau51]. commentary [DP97]. Comments [Pea24, Ree22]. Commission [BBD^{+}46]. Comparisons [Mit86]. components [FP31]. Composite [Fri39a]. Comprehensive [GJ23]. Concepts [Mit86]. conceptual [Fri71, Fri09b]. Concerning [HS39]. conditions [Per39]. Conference [KLR13]. conservation [Fri59j]. Conserved [Fri59b, Fri57e]. consideration [CSW97]. Considerations [Fri59c]. Construction [FP40c, FP64, FP65, Arn03]. Contemporary [Hen80, Fri62c]. Continuity [Mit86]. Contributions [FPLR59a]. control [FCJ60, BGR61, BGR61]. conventions [DWS79]. Copenhagen [Fri79a, Sch00]. correspondence [ALJ82]. Cosmology [Rei72, Gar63a]. counter [FF46]. counters [Fri50d]. Counting [Fri55e]. Countries [Arm66b]. course [Rot54]. Crawford [Bar05]. creation [Stu13]. Creator [Fri73a]. crimes [Wes04]. Critical [Pea24, Ber02, DWS79, McC24, McC25]. Crystal [GT15, Fri33a, FS33d, For17]. Crystals [FO59]. culture [Ros98]. Curies [Kae48]. curious [FL06]. current [DWS79]. currents [FM28a]. Curves [Fri39a]. cyclotron [Fri50b]. Cyclotrons [Row09].

dagegen [CSW97]. Danish [BR64, Fri64g, RRK^{+}64]. Dardo [Bar05]. dark [Fri75b]. David [BGR61, Smo07, Sta08]. Davis [Fri73a]. dawn [Rif06]. Day [Lau51]. Days [Bar87, WM15]. Decay [FTW47, Fri39a]. decision [Fri60b]. dedicated [FPLR59b, Lin60]. defeat [Ols63]. Defining [Mla98]. Definitions [KED^{+}60]. deflection [FS33e]. DeGregori [Mit86]. delay [FF46]. Delayed [SFFdH45]. Delhi [Fri57b]. Dennis [Smo07]. deposited [ALJ82]. dergleichen [Fri28]. Description [Boh61]. Designer [Fri55d]. Detection [FO59]. deterrence [GA74a, GA74b]. Develop [WHF^{+}96]. Development [Mit86]. device [Fri28, FM28a]. did [Ols63]. Dies [Fri68c]. Different [Fri30a, FHS31]. diffraction [Fri32, Fri33a, FS33d, FS31]. Diffusion [Arm66a]. direction [FS33c]. directional [FS33a]. Discoloration [Fri26]. discovered [Fri57d]. discoveries [Lig05]. discovers [Ell08, Trö09]. Discovery [FW67, Gar63b, Gra64, Hah58, Mei62, Sim07, Fri56d, FW68, FW85, GA71, Kae48, Mül68, Sim14]. Discussions [Bas71, Bas09]. Disintegration [Boh39, MF39a, MF68, MF69]. dispute [Fri66a]. Distinguished [Ano73, Fri73d]. Distinguishing [Fri74c]. distribution [FP31]. Division [Fri39c]. do [Fri53b]. documentary [GA71]. Doubtful [Kae39]. down [FvK37e]. dreamers [Poh62]. Drehimpulsbilanz [Fri30b]. drei [Ell08]. Driven [Fri54a, Fri56f]. Drop [Ano94, Stu94]. dualism [Fri61d]. durch [Fri26, Fri30a]. during [Rot54].

Each [Ano39]. Early [Bar87, Fri77b, Rot00]. earth [HS39]. economic [McC01]. ed [Sta08]. Edited [For17, Lin60, Ano64, BGR61, Fir64, Fri64d, Fri67a, Smo07]. Editor [Bar05]. eds [Fri74a]. education [Gar63a]. Effect [Fri27]. effects [PF40, Fri57b]. Einstein [Fri73a, Fri33b, Fri79c, Nef05, Nef06, Nef07, Smo07, Sta08, Fri69, Smo07]. Einsteinschen [Fri33b]. Einstellung [FS33c]. Einstrahlung [FS33a]. Electron [Ano39, Fri32]. Elektronenbeugung [Fri32]. Element [Nod34]. Elementarteilchen [FF63]. elementary [Fri59h, Fri61f, FF63]. Elements [FvK37a, Sim07, Hab09]. Elisabeth [Bar05, Bar05]. Emilio [Seg93]. emission [Fri30b]. Empires [Smo07]. Encyclopaedia [DWS77]. Encyclopedia [Bar05]. energetic [Fri59f]. Energieerzeugung [Fri57c]. energies [Fri50a]. Energy [Ano39, BBD^{+}46, Gow64, Mei62, Seg85, Fri54h, Fri54d, Fri55f, FCJ60, GA74a, GA74b, Rot54]. enjoys [Pei79]. Enrico [Fri55d, Fri55b]. Entdecker [Ell08, Trö09]. entstehenden [HS39]. Epoch [Ano39]. EPS [HJP17]. Equilibrium [FvK37f]. Erdalkalimetalle [HS39]. erhielt [CSW97]. erinnere [Fri81]. Ernests [Oli66a, Oli66b, Oli85]. Ernst [Lee07]. Errata [Ano94]. Erwin [Pra93]. Essay [Bar05]. Essays [Bas71, Bas09, Boh61, FPLR59b, Lin60]. Essence [Sta08]. establishment [Ols63]. Estimate [Pea24]. etc [FM28a]. éventuelles [Per39]. Evidence [Fri39c]. Excellence [Bar05]. existence [HS39]. expect [Fri53b]. Expelled [MP01]. Experimental [Fri79a, Fri33b]. experimentalist [Fri71, Fri09b]. Experimenteller [Fri33b]. Experiments [FvK38, EFS31]. expert [Poh62]. Explaining [Bro47, Bro47, GHJ47]. Explanation [Sta08]. Exploring [Fri60c]. explosion [Fri70a]. explosions [Fri57b].

F [Fri73b, Lin60, Lip59]. F. [Arm66b]. F.R.S [AP77]. F.R.S. [Fri75a]. facsimiles [Bey49]. Fact [Pei86]. Failed [WHF^{+}96]. Family [Fri55d]. Farm [McC24]. Farrar [Sta08]. fast [DFS70, PF40]. Fear [WM15]. Feasibility [Fri54a, Fri56f]. February [MF69, Rot54, Arn03]. Feld [Fri74a]. Fermi [Fri55d, Fri55b, Bye02]. Ferromagnetism [Bet33]. Ferromagnetismus [Bet33]. Feynman [FL06]. Field [FvK37b, KED^{+}60]. Fifty [Kae48]. First [Fri55d, Sim96, WM15, Fri70a, Fri77c, SR92, Stu18]. Fission [Bar87, FW67, Gar63b, Gra64, Hah58, MF39c, Pea24, Sim96, Sim07, Sim12, Stu85, Stu94, Ell08, Fri57d, Fri57f, FW68, Fri74e, FW85, GA71, Kra14, MF39b, Mül68, Pei86, Sim14, Stu13, Trö09, Whe09, Ano94, CSW97]. flash [Gar63c]. Fled [MP99]. fluctuations [FL54]. Fluorine [Fei35]. focussing [Fri67b]. Force [Ano39]. Forgetting [Sim12]. fortalt [BR64, RRK^{+}64]. Found [Sim96]. Foundations [Bey49]. Four [Boh61, Trö09]. Fred [Smo07]. Free [FvK37c, FvK37d, FvK38]. Freed [Ano39]. Freeman [Bar05]. French [Fri55a, Gol87, Per39]. Friedman [Bar05]. Friedr [Lip59]. friends [BR64, RRK^{+}64, Roz67]. Frisch [ALJ82, BGR61, But64, CSW97, Fel79, Hen80, KED^{+}60, Lin60, Lip59, Pei79, Pei81, Smo07, Sta08, Ano64, Ano73, Arn03, BW80, Bro47, CSW97, Dit80, Fes54, Fir64, FP40a, FP97, GHJ47, Gar63b, GJ23, Hal61, Kuh63, McC25, Mil09, Pea24, Pei80, PR04, Pra93, Ree22, Ree24b, Sei80, Sie80, Sim96, Smi61, Stu13, Wei79, Wil80]. Frischtak [Mit86]. Fritz [CSW97, CSW97]. FRS [ALJ82]. Fundamental [Bha39, Fri60d, Fri61h, Bey49, Fri63b, Fri64h]. fusion [Rei72].

G [Fri62b, Pra93]. Galison [Smo07]. galore [Fri67c]. game [Ree15a]. game-changer [Ree15a]. Gamow [Pra93, Rei72, Smi61]. Gebieten [FHS31]. Generation [Fri57c]. genius [Gar63c]. geopolitical [Ree15a]. George [Bar05, Fri55d, Fri75a, Pra93, Smi61, Rei72]. Gerät [Fri28]. German [Smo07, Bet33, CSW97, Ell08, EFS31, EFS32, Fri26, Fri27, Fri28, FM28a, FM28b, Fri30b, Fri30a, FS31, FP31, FHS31, Fri32, FS32, Fri33a, FS33d, Fri33b, FS33c, FS33a, FS33e, FFPE35, Fri57c, Fri59c, FPLR59a, FF63, Fri81, HS39, Nef05, Nef06, Nod34, Ols63, Ros98, Sim12, Trö09, Wal03].

Germany [Lin60, MP99, WHF^{+}96]. Gertrude [Fri74a]. Giant [Fri53b]. Gift [MP99, MP01]. Giroux [Sta08]. given [Rot54, Rot00]. Glazer [For17]. global [Ree15a]. Gold [FHM36]. Got [Sim96, Ols63]. Great [Fri79c, Fri60b, Kae48, Lig05, Ols63]. greatest [McC01, Wes04]. group [RRK^{+}64, Rot00]. groups [Fri42b]. Grundlagen [KED^{+}60]. guide [Fri47b, Fri47c].

H [Bar05, BGR61, Hen80, Pra93]. Hadley [BGR61]. Hahn [CSW97, Lin60, Sto89, Wal03, CSW97, Fri51d, FPLR59b, Fri68a, Hof01, Sim12, Wal03, Wal06]. Hahn/Lise [Sto89]. Hailed [Ano39]. Hall [McC24]. Halperin [BGR61]. Hamburg [Fri79a]. handbook [Fri58a, Fri58b]. hans [BR64]. Hardy [Pra93]. Hargittai [Bar05]. Hart [Fri73a]. Hart-Davis [Fri73a]. Health [KED^{+}60]. Heavy [Boh39, Fri39c]. Hecht [Bro47, GHJ47]. Heisenberg [KED^{+}60, Sch00, Ber02, McC24, Ree24a, Ros98]. held [AP77]. hemispheres [FP43]. Herewith [Ols63]. Hewlett [Fri62b]. Hg [FP31, FP31]. Hg-Triplett [FP31]. hidden [Far13]. High [Fri55f, Fri50a]. High-energy [Fri55f]. Highfield [Smo07]. Hiroshima [Lau51]. Historic [HJP17]. Historical [Seg85, Bar05]. Histories [Pei97, Wei79]. History [Hen80, WP85, Arn03, Far13, GA71, KLR13, Wes04]. Hitler [Dei96, MP99, MP01]. Hoffman [Fri73a]. House [HJP17]. HQ [KLR13]. HQ-3 [KLR13]. Human [Mit86]. Humanitarian [Fri79c]. Hundred [Bar05]. Hydrogen [Fri57c, FS33e, Fri54e, Fri54g]. hydrogenous [FvK37e].

ich [Fri81]. Ida [Hab09]. Ignorance [DWS77]. ihr [CSW97]. II [Fri33a, FS33c, FS33a, Fri55a, Fri62a, Oli66b, Sim12]. illus [Sta08]. im [FP31]. images [Fri28]. impact [Fri62c]. Impossible [Fri78]. Impulse [GJ23]. including [McC25]. Independence [GA74a, GA74b]. index [Bar05, Sta08]. Indian [Fri57b]. Individuality [Fri54f]. Induced [Fei35, Fri34]. Information [Fri57b]. injustice [CSW96]. inneres [Fri32]. Innocence [Stu18]. Institute [Fri50a]. Intensitätsverteilung [FP31]. intensities [FM28a]. intensity [FP31]. interactions [Fri59j]. Interessen [Fri64g]. interest [Fri67b, Fri64g]. internal [Fri32]. International [KLR13, Mit86]. Interpretation [Ano94, Stu94, Stu13]. Interscience [Lin60]. Interview [Kuh63]. Introduced [Bha39]. Introduction [Fri61g, KED^{+}60, Fri64a]. introductory [Boh61]. Ionisationsvorgange [FM28a]. ionization [FM28a]. Iron [FvK37b]. irradiation [FS33a, HS39, SFFdH45]. Isaacson [Smo07, Sta08]. ISBN [For17]. ISBN-13 [For17]. Isotope [Fri42b]. Issues [BGR61]. István [Bar05]. Italian [FS33b].

J [Fri69]. Jahrhunderts [KED^{+}60, FPLR59a, Lip59]. Jahrhunderts/Bellman [KED^{+}60]. January [Rot54]. Jelley [Fri59e]. Jerome [Smo07]. Jocular [Fri56a]. John [Bro47, GHJ47]. journals [Bey49]. Journey [FR13]. July [AP77, KLR13]. June [KLR13, Lee07]. Jürgen [Smo07, Sta08].

Kathodenstrahlen [Fri27, Fri26]. Keeping [Bal14]. Kernspaltung [CSW97, Ell08, Trö09]. key [Arn00]. kick [FFS51]. kick-sorter [FFS51]. kohärente [Fri59c]. Kowarski [Fel79, Fri79d]. kreds [BR64, RRK^{+}64]. Kristallspaltflächen [Fri33a, FS33d]. Kungälv [HJP17]. Kurian [Bar05].

Lady [For17, GT15]. Landé [Fri61d]. langsamen [Fri27]. Laue [Lin60, FPLR59b]. Laura [Fri55d]. Laureates [Bar05, McC01]. Laves [Lin60, Lip59]. Lawrence [For17, GT15]. Laws [Fri60a, Fri60g, Fri66a]. Layers [Fri30a]. Laymen [Smi61]. Leaders [Lau51, Ols63]. Lecture [Mar54]. Lecturer [Ano73]. Lectures [Fri47a, Fri59d, Rot54, SR92]. Legacy [Kel04]. Length [Bha39]. Leo [Fri74a]. Letter [Fri56e]. Leuchtdauer [FP31]. Lew [Fel79, Fri79d]. Liberate [Ano39]. Library [ALJ82]. Lichteinheit [FM28b]. Lichtemissionsvorgängen [Fri30b]. Lichtintensitaten [FM28a]. Life [Arm66b, Coc46, Fri55d, Fri79c, Mar54, Pra93, Smo07, Sta08, BR64, Fre86, FP31, Fri67a, RRK^{+}64, Roz67]. Light [FM28b, FvK37a, FP43, FM28a, Fri30b]. like [Fri28]. limitations [Fri50b]. lines [FF46]. Liquid [Ano94, Stu94]. Liquid-Drop [Ano94, Stu94]. Lise [CSW97, Lin60, Sto89, CSW97, FPLR59b, Fri68c, Fri70b, Fri73d, Fri74d, Gar63b, Rif06]. Little [Fri79e, Hen80, Pei79, Pra93, Sei80, Sie80, Wil80]. liv [BR64, RRK^{+}64]. Liverpool [Row09]. Lives [Smo07]. London [Fri50a, Fri55d, Fri73a, Fri73b, Fri74a, Fri79a, Rot54]. Lord [Smo07]. Love [Smo07]. Ltd [But64, Fri55d]. LXIII [FFS51]. Lying [DWS79].

M [For17, Fri60b, Fri62c]. MacGibbon [Fri73a]. machine [DFS70]. Magnetic [EFS33, FvK37b, FvK37c, FvK37d, FvK38, FS33e, Fri63a]. magnetische [FS33e]. magnetism [Fri59l]. Magnetized [FvK37b]. main [Fri74c]. Making [Ano39, Kel04]. man [Fri65c]. Mandl [KED^{+}60]. Manhattan [Kel04, Ree15a, Sza92a, Sza92b, Wes04]. Maps [Smo07]. Marc [Bar05]. March [Rot00]. mark [AP77]. Marks [Sch00]. Marley [KED^{+}60]. Mass [Pea24, Ber02, McC24]. Materiestrahlen [FS31]. Mathematical [GJ23]. Mathematician [Pra93]. mathematics [Gar63a]. Matrix [KED^{+}60]. Matter [FTW47, Fri72b, Fri73e, Pra93, FS31, Fri56b, Fri57g, Fri64e]. matters [Rei72]. Mauro [Bar05]. Max [Lin60, FPLR59b]. Means [Fri63b, Fri42b]. Measurement [KED^{+}60]. Measures [Mit86]. Measuring [FvK37c, Fri59g, Fri72a, Fri73c, DFS70, Fri28]. mechanical [FFS51]. Mechanism [FW67, Whe09]. medarbejdere [BR64, RRK^{+}64]. Medicine [Bar05]. Meet [Fri47b, Fri47c, Bro47, GHJ47]. Mehlin [KED^{+}60]. meiner [Fri81]. Meitner [CSW97, Fri70b, Lin60, CSW97, FPLR59b, Fri68c, Fri73d, Fri74d, Gar63b, Rif06, Sim96, Sto89, Stu13]. Mem [Fri55b]. Memo [BBD^{+}46]. Memorandum [FP40a, FP40b, FP97, FP04, Ree22, Ree24b, Arn03, Ber11, McC25]. Memorial [Mar54, Rei72]. Meson [Bha39, Fri48a]. Mesotron [Bha39]. Mesotrons [FTW47]. Messages [FFPE35, FHS31]. Metalle [Fri32]. metals [Fri32, HS39]. Method [FvK37c]. Mica [AFL39]. mich [Fri81]. Mind [Pra93, Seg93]. minds [McC01]. Ministry [Fri57b]. missing [Hab09]. Mitteilungen [FHS31, FFPE35]. mittels [HS39]. mittlere [FP31]. Model [Ano94, GJ23, Stu94]. Modern [Smi61, Fri47b, Fri47c, Fri53a, Fri54f, Fri65e, Fri66b, McC01]. modulation [FL54]. Molecular [Fri59a, Fri65a, EFS31, EFS32, FS32, Fri33a, FS33d]. molecules [FS33e]. Molekularstrahlen [Fri33a, FS33d, EFS31, EFS32, FS32]. Moment [EFS33, FS33e, FvK37c, FvK37d]. momentum [Fri30b]. Monday [Rot00]. Monochromasierung [EFS32]. monochromatic [EFS31]. monochromatischen [EFS31]. Monochromatization [EFS32]. Moore [Fri67a]. Morgan [KED^{+}60]. Mortal [Wes04]. Morton [BGR61]. motion [Seg93]. Move [GHJ47]. Mr. [Pra93]. multiplicative [Fri48b]. My [Fri55d, Fri81, Rot00]. Mystery [FR13].

N.Y [Bar05]. Nachweis [Fri33b, HS39]. names [Kra14]. Nathan [Mit86]. Nation [BGR61]. national [Bal14]. Nature [Boh61, Fri60a, Fri60g, Fri72b, Fri73e, FCJ60, Fri64e, Bal14, MF69, Sim96]. Naturwissenschaft [KED^{+}60]. Naturwissenschaft/Heisenberg [KED^{+}60]. Nazi [MP99, MP01, Ros98, WHF^{+}96]. Nazism [Ols63]. Neffe [Smo07, Sta08]. Negative [FTW47]. neptunium [Nod34]. Neutron [Fri39c, Kae39, Fri42a, HS39, MF39b]. Neutronen [HS39]. Neutrons [FS35, FP36, FHM36, FvK37a, FvK37b, FvK37c, FvK37d, FvK37f, FvK38, MF39a, MF68, MF69, FvK37e, Fri37, PF40, SFFdH45]. News [Stu85, Rot54]. nicht [CSW97]. Niels [BR64, Fre86, Fri67a, RRK^{+}64, Roz67, Fri67a]. Nobel [Bar05, Bar05, CSW96, CSW97, McC01]. Nobelpreis [CSW97]. Noddack [Hab09]. Nominators [Bar05]. Nominees [Bar05]. non [Fri59j]. non-conservation [Fri59j]. Note [FF46]. notes [Bar05]. Nuclear [Ano64, Ano73, Ano94, But64, Fes54, Fir64, FP40c, Fri50c, FCJ60, FP64, Fri64c, FP65, Fri68c, Fri73d, Gra64, MF39a, Mei62, MF68, MF69, Mla98, Ros72, Seg85, Sim07, Sim12, Stu94, Stu18, And73, Arn00, Arn03, Bey49, Ell08, Fri46, Fri47a, Fri57d, Fri58a, Fri58b, Fri61a, Fri64b, Fri70a, Fri74e, GA71, Kae48, Ree15a, Rif06, Sim14, Stu13, Trö09, Wil75, Hen80, Stu79, Fri57b]. Nuclear-Fission [Sim12]. Nuclear-Physics [Fri68c]. Nuclei [Boh39, Fri39c, Fri64b, Fri79a]. Nucleus [FR13, MF39c, Fri59f, Fri59k, Fri64g, Fri67b, Smi61]. nur [CSW97].

O [Ano64, Bro47, But64, Fir64, Fri62b, GHJ47, Lin60, Lip59, Mil09, Sie80]. O. [Gar63b, Pei80]. Obituary [Fri75a, FM79, Fri79d, Pei80, Dit80]. Observation [Fri62d]. observing [Fri63b]. occasion [AP77, FPLR59b, Lin60]. October [Pei81]. omkring [Fri64g]. once [Ols63]. once-great [Ols63]. One [Bar05]. only [Ano39, CSW97]. Open [Fri68b]. Opening [Sch00]. Optics [KED^{+}60]. Optics/Mandl [KED^{+}60]. Oral [Wei79]. ordinary [PF40]. Origin [Ano94, Stu94, Fri74e]. originally [Bey49]. other [Rei72, Smo07]. Otto [ALJ82, CSW97, Fel79, Fes54, Hal61, Hen80, Lin60, Pei79, PR04, Pra93, Sei80, Smi61, Wil80, Ano73, BW80, CSW97, FPLR59b, Fri68a, Hof01, Kuh63, Pei81, SBRT^{+}16a, SBRT^{+}16b, SBRT^{+}16c, SBRT^{+}16d, Sim12, Sto89, Wal03, Wal06, Wei79]. OUP [For17]. our [Wil75]. Overbye [Smo07]. Oxford [AP77, Bar05, But64, Fri59d]. Oxford/New [Bar05].

P [Lin60, Lip59]. Paget [Fri75a]. Pais [Smo07]. Paneth [Lin60, Lip59]. Paperback [Pra93]. Papers [Smo07, ALJ82, DP97, Fri74a]. Parity [Fri57e, Fri59j, Fri60f, Fri59b]. Part [FP40a, FP65, Fri61b, Fri62a, FP64]. particle [Fri77c]. Particles [Fri51a, Fri50a, Fri51c, Fri55e, Fri56b, Fri59h, Fri59i, Fri60d, Fri60e, Fri61f, Fri61h, FF63, Fri63b, Fri64h, Fri65e, Fri66b]. Patience [For17]. Paul [Smo07]. Peaceful [Fri57f]. Peierls [AP77, Lee07, Arn03, DP97, FP40a, FP97, McC25, Pea24, Ree22, Ree24b]. penetration [Wes04]. Penguin [Gar63a]. People [Bro47]. Pergamon [But64]. period [Coc46]. personal [Coc46]. Perspective [Seg85]. Perspectives [Kel04]. Peter [Smo07]. phenomena [FP39, Fri39b, Fri40, Fri41, Fri43, OPM38]. philosophical [Fri62c]. Philosophie [KED^{+}60]. Philosophie/Frisch [KED^{+}60]. Phosphorus [Fri34]. photographs [Fri28, FO60]. photon [Fri65b, Fri09a, Kra14, Mil09]. Physical [Fri39c, FFPE35]. physicist [ALJ82, Pei79, PR04, Rot00]. Physicists [Fri81]. Physics [Bar05, Fri50c, Fri56a, Fri59b, Fri64c, Fri68c, Fri81, Hal61, Hen80, KED^{+}60, Lin60, Mla98, Pei91, RW64, Ree22, Smi61, Stu18, AP77, Bal14, Bey49, Fri46, Fri47a, Fri47b, Fri47c, Fri53a, Fri54f, Fri55a, FPLR59a, FPLR59b, Fri61c, Fri61a, FF63, Fri65e, Fri66b, Fri67d, Fri74f, Gar63a, Kae48, KLR13, Kra14, Ree15a, Stu79, WP85, Bar05, Fri50a, Ano64, But64, Fes54, Fir64, Fri59d, Fri61d, Fri62c, Fri64d, Fri73b]. Physics/Churchman [KED^{+}60]. Physik [FPLR59a, Fri81, KED^{+}60, Lip59, FF63]. Physikalische [FFPE35]. Physiker [Fri81]. physique [Fri55a]. Pile [FL54, Fri55d]. piles [FL54]. Pioneer [Ano73, Fri68c, Fri73d]. pioneers [Wil75, Gol87]. Pionniers [Gol87]. plasma [Kra14]. plates [Fri55d]. Poincaré [Smo07]. point [Fri71, Fri09b]. points [Fri59d]. Poland [Rot00]. Politics [Bar05, Sim12, Far13, Sim14, Smo07, Sta08]. Poor [Fri65c]. popular [Fri47b, Fri47c]. Population [Bar05]. Possibilities [Kae39]. possible [Arn03, Per39]. Potential [Fri32]. Power [Fri54a, Kae39, Fri56f]. Pp [But64, Fri55d, Fri57b, Fri61d, Lip59, Bar05, Fri73a, Fri73b, Fri74a, Lin60, Pei79, Sta08]. Prazisionsgerat [FM28a]. precision [FM28a]. Predictions [McC25]. Preis [CSW97]. Presenting [Fri74f]. Press [Bar05, But64, Fri61d, Fri73b, Fri74a, Pei79, Sta08]. price [CSW97]. primer [SR92, McC24]. Princeton [Sta08]. Principles [KED^{+}60]. Private [Smo07]. Prize [Bar05, CSW97]. Prizes [Bar05, Bar05]. Probing [Fri59k]. problem [Fri71, Fri09b]. Probleme [Fri57c]. Problems [Fri57c]. Proceedings [Hen80, AP77, Stu79]. Process [Ano39, FPSS32, Gar63b]. processes [FM28a, Fri30b, Fri48b]. Produces [Ano39]. Products [MF39c, MF39b]. Prof. [Fri55b]. Professor [AP77]. Program [BGR61]. Progress [Fri50c, Fri61a, Ano64, But64, Fes54, Fir64]. Project [Arm66a, Ros98, Kel04, Ree15a, Sim12, Sza92a, Sza92b, Wes04]. Prometheus [Bar05, Fri79b]. proof [Fri33b]. Propagation [GJ23]. Properties [FvK38, FP40b, FP40d, FP04]. Prophet [Arm66b]. Proton [EFS33, FS33e, Fri55c, Fri56c, Fri56d, Fri63a]. Protons [FS33e]. prove [Fri79f]. publication [Bal14]. Publishers [Lin60]. publishing [Bal14]. pulse [FFS51, FF46].

Quantisation [FPSS32]. Quantization [FS33b, FS33c, FS33a]. Quantizzazione [FS33b]. Quantum [Bas71, Bas09, KED^{+}60, Fri62d, Fri71, Fri09b, KLR13]. quarks [Fri77a]. quasars [Fri77a]. quest [Fri57g]. questions [And73].

R [Ano64, Bro47, But64, Fes54, Fir64, Fri55b, Fri62b, GHJ47, Gar63b, Hal61, Lin60, Lip59, Mil09, Mit86, Pei79, Pei80, Sei80, Sie80, Smi61, Wil80]. Race [Lau51, Smo07, WM15, Bal14]. Racism [Smo07]. radar [Fri48a]. radiation [Fri33b, Fri59e]. radii [McC25]. Radioactive [FP40b, FP40d, FP04]. radioactivité [Fri55a]. Radioactivity [Fei35, Fri34, FP39, Fri39b, Fri40, Fri41, Fri43, OPM38, Fri55a, Fri74c]. Radium [Kae48]. Rational [FM28b]. Rationellen [FM28b]. ray [Fri42b]. Rays [Fri30a, Kae39, Fri26, Fri27, FS31]. Re [Pra93]. Re-Reading [Pra93]. Reaction [FP40c, FP64, FP65, MF39a, MF68, MF69, And73, Fri77b]. reactions [Fri64b]. reactor [McC24]. reactors [McC25]. Reading [Pra93]. Realisierung [FM28b]. realization [FM28b]. realm [Kae48]. Rebel [Fri73a]. Recalling [Arn00]. Recalls [Lau51]. Received [Ano64, CSW97]. recoil [Fri33b]. recording [Fri28, Fri74c, Fri74e]. Reduction [BGR61]. reflection [FS32, Fri33a, FS33d]. Reflexion [FS32, Fri33a, FS33d]. Refugee [Seg85]. Regime [MP01]. Registering [FM28a]. Registrieraufnahmen [Fri28]. Registrierendes [FM28a]. relatif [Per39]. Relativist [Fri79c]. relativity [Fri61b, Fri62a, Fri64a]. Remember [Fri79e, Fri81, Hen80, Pei79, Pra93, Sei80, Sie80, Wil80]. Remembering [Kel04]. Reminiscence [Hen80]. Reminiscences [Bar87, Fri51b, Ros72, Coc46, Wil75]. Report [ALJ82, Kae45]. reported [Bey49]. Repression [Wal06, Wal03]. Reprinted [MF69]. research [Gar63a]. Researches [FS33b]. Responsibility [Wal06, Wal03]. Rest [Sim96]. resulting [HS39]. retirement [AP77]. Retrospect [Hen80, Stu79]. Revealed [Kae45]. Review [Bar05, But64, Fes54, Fir64, For17, Fri50a, Fri51d, Fri55d, Fri57b, Fri59e, Fri59d, Fri60b, Fri61e, Fri61d, Fri62b, Fri62c, Fri64a, Fri64e, Fri64d, Fri69, Fri73a, Fri73b, Fri74a, Fri77a, GHJ47, Hal61, Lin60, Lip59, Mit86, Pra93, Sei80, Sie80, Smi61, Wil80, Ols63, Pei79]. Reviews [Fri57a, Hen80, Sta08]. Revisiting [Ree24b]. Revolution [Kae48]. revolutions [And98]. Ricerche [FS33b]. Richtungsquantelung [FS33c, FS33a]. Right [Mei62]. Rivals [Gol90]. Road [Bar05]. Roads [Mei62]. Robert [ALJ82, Bar05, BW80, Fel79, Pei81, Smo07, Sta08, PR04]. Robin [Fri61e]. rock [Fri26, Fri27]. Rodger [Smo07]. Roger [Hen80, Smo07]. role [Arn00]. Romance [Smo07]. Rosbaud [Lin60, Lip59]. Rosenberg [Mit86]. Rosser [Fri64a]. Rowe [Smo07, Sta08]. Rozental [Fri67a]. Rudolf [AP77, Lee07]. Rudolph [AP77]. Runcorn [Fri64d]. Ruth [Fri67a]. Rutherford [Coc46, Mar54].

S [Fri64d, Fri67a]. S. [Fri55b]. Safety [BGR61]. salt [Fri26, Fri27]. samler [Fri64g]. samples [Fri42b]. scattered [Ols63]. Schelling [BGR61]. Schichten [Fri30a]. Schrödinger [Pra93]. Schulmann [Smo07, Sta08]. Schuster [Sta08]. schwache [FM28a]. Schwachung [Fri30a]. Science [Bar05, HJ60, Hen80, Kae39, Sim14, Smo07, Far13, Gar63a, Lig05, Ree15b, Bar05]. sciences [Rei07]. Scientific [Fri74a, Smo07, And98, Bey49, DP97, Ols63]. Scientist [Fri68b, Kae45, Ols63, Fri56e]. Scientists [Bar05, Fri54g, MP99, MP01, Seg85, Sza92a, Sza92b, Bar05]. Scintillation [Fri50d]. scrutiny [DWS79]. Search [Sim07]. Second [Stu18]. Secret [Kae45]. seen [Roz67]. Segrè [Seg93]. sehr [FM28a]. Selected [DP97]. Selective [FHM36, Fri37]. Selig [Bro47, GHJ47]. semi [DFS70, FO60]. semi-automatic [DFS70, FO60]. September [Lee07, Pei81]. Serber [McC24]. setting [FS33c]. shaped [McC01]. Shelley [Smo07, Sta08]. Shooting [Fri75b]. short [SFFdH45]. sig [Fri64g]. Sign [FvK37d]. Simon [Sta08, Arm66b]. Sir [AP77, For17, Fri75a, GT15]. Sites [HJP17]. sixties [Fri64d]. Sixty [FR13]. size [FFS51]. Sketches [Pra93]. slits [Fri33a, FS33d]. Slow [FS35, FP36, FvK37a, Fri27, Fri37]. slowing [FvK37e]. slowing-down [FvK37e]. SMP [Fri65c]. Smyth [Kae45]. snow [Fri75c]. Sodium [Fri34]. Sohn [Lin60, Lip59]. Some [FvK38, Ols63]. Somebody [Fri74b]. sorter [FFS51]. Sound [Fri74c, Fri74e]. Source [Kae39, Lei68, Fri42a]. Soviet [Wes04]. Space [FPSS32, Gar63a]. sparrows [Fri75b]. Spatial [FS33b]. Spaziale [FS33b]. spectral [Fri28]. specular [FS32, Fri33a, FS33d]. speculation [Pei86]. speed [Bal14]. Spektralphotographien [Fri28]. spiegelnde [FS32]. spiegelnden [FS33d]. Split [Ano39]. Splits [Kae39]. Stations [Fri54a, Fri56f]. Statistical [Fri39a, FL54]. Statistics [Fri48b, Fri79f]. Steinsalz [Fri26, Fri27]. steps [Fri77b]. Sterns [SBRT^{+}16a, SBRT^{+}16b, SBRT^{+}16c, SBRT^{+}16d]. still [Kae48]. Stockholm [Bar05]. Story [Kae45, MP01, Ree15a, Bro47, GHJ47]. Strahlen [Fri30a]. Strahlungsrückstoßes [Fri33b]. strange [Fri60e]. Strangeness [Fri60f]. Straßmann [CSW97, CSW97]. strategies [Bal14]. Strategy [BGR61]. Straus [Sta08]. Strome [FM28a]. Studies [Bar05, Bey49]. study [Ros98]. Stuewer [Hen80]. sub [FP39, Fri41, Fri43, Fri60c, Fri77c, OPM38]. sub-atomic [FP39, Fri41, Fri43, Fri60c, Fri77c, OPM38]. subatomic [Fri39b, Fri40]. substances [FvK37e]. Subtle [Smo07]. Sulla [FS33b]. Sun [Fri74b]. Super [FP40c, FP40b, FP40d, FP64, FP65, FP04]. Super-bomb [FP40c, FP40b, FP40d, FP64, FP65, FP04]. Surprises [Pei91]. Survey [HJ60, Boh61, Gar63a, Rot54]. sustained [And73]. Sweden [HJP17]. Sweepnik [DFS70]. Switch [Fri74b]. Symmetrical [Fri60a, Fri60g]. Symmetry [Fri66a]. Symposium [Hen80, AP77, Stu79, Sch00]. synchrotron [Fri55f]. Szilard [Fri74a, Bye02].

T [BGR61, Fri74a]. Take [Fri65b, Fri09a, Mil09]. taken [CSW97]. tale [CSW96]. talk [Rot00]. Tau [Fri30a, Fri30a]. Tau-Schichten [Fri30a]. Taylor [Smo07]. Technology [Mit86]. Temperature [FvK37f]. theft [Wes04]. their [Fri57b, FPLR59b, Lin60]. them [Fri53b, Ols63]. Theoretical [Pei91, AP77]. Theorie [Bet33]. Theories [KED^{+}60]. Theory [Bas71, Bas09, Bha39, Boh61, KED^{+}60, Mit86, Sim96, Fri48a, Fri64a, Fri71, Fri09b, McC24, Bet33]. Theory/Marley [KED^{+}60]. there [Ols63]. Thin [AFL39]. things [Fri78]. Third [KLR13]. thirteen [Bey49]. Thomas [Bar05, BGR61, Mit86, Fri77a]. Thomson [For17, Fri75a]. thorium [MF39b]. thought [McC01]. Thousand [WM15]. Three [And73, Ell08]. Time [Fri61b, Fri62a, Fri81, Wil75, Smo07]. Times [Fri79c]. Tizard [Cla65]. today [Fri61c, Hal61, Smi61]. Tokyo [Bar05]. told [BR64, RRK^{+}64]. Tompkins [Pra93]. Tools [Fri64c, Fri46]. track [DFS70]. track-measuring [DFS70]. Trackdown [Ols63]. Tracks [Fri67c, Fri72a, Fri73c]. Traditions [KLR13]. Trans [Sta08]. Transfer [Mit86]. transformations [KLR13]. translated [Smo07]. transmutation [Per39]. Trends [FPLR59b, Lin60]. Trinity [ALJ82]. triplet [FP31]. Triplett [FP31]. Triplettkomponenten [FP31]. True [MP01]. truths [DWS79]. Turned [Fri74b]. Turning [Fri59d]. twelve [Wil75]. Twentieth [Bar05, Lig05]. Twentieth-Century [Bar05, Lig05]. Twist [Fri59b]. Two [Arm66b, Lau51, FP43, Oli66a, Oli66b, Oli85]. Type [MF39a, MF68, MF69]. types [Fri74c].

U [SFFdH45]. Uddmanska [HJP17]. unbounded [Kae48]. Unit [FM28b]. unity [Fri61d]. Universal [Bar05]. Universe [Sta08, Smo07]. University [Bar05, Fri61d, Pei79, Sta08, Lau51, Rot54]. Unravelling [FR13]. Unwin [Fri55d]. upon [FvK37b]. Uranium [Ano39, FP40c, FP64, FP65, Kae39, MF39a, MF39c, MF68, MF69, Pea24, Fri42b, HS39, MF39b, PF40, Per39]. Urans [HS39]. use [FF46, FCJ60]. Uses [Ano39, Fri57f]. usw [FM28a].

V [Fri59e, Fri64a, Fri73b, But64]. Vast [Ano39]. Velocity [FS35]. venner [BR64, RRK^{+}64]. Verantwortung [Wal03]. Verdrängung [Wal03]. Verfärbung [Fri26]. Verhalten [HS39]. Veröffentlichungen [SBRT^{+}16a, SBRT^{+}16b, SBRT^{+}16c, SBRT^{+}16d]. verschiedenen [FHS31]. verschiedener [Fri30a]. Versuche [EFS31]. very [FM28a]. vi [Bar05]. vier [Trö09]. view [Fri71, Fri09b]. Vieweg [Lin60, Lip59]. views [Rot54]. viii [Bar05]. virke [BR64, RRK^{+}64]. vision [Fri69]. Void [GHJ47]. Volt [Ano39]. Volts [Ano39]. Volume [Ano64, Fir64, But64, Rei72].

W [Bar05, Bro47, GHJ47]. walk [Fri75c]. Walter [Smo07, Sta08]. Wandlungen [KED^{+}60]. War [Sim12, Far13]. Warburg [FM28b]. Warburgschen [FM28b]. Wars [Stu18]. wartime [CSW96]. Warum [CSW97]. Wasserstoff [Fri57c]. Wasserstoffmolekülen [FS33e]. Wavelengths [Fri30a]. waves [EFS31, EFS32]. Way [Fri68b]. weak [FM28a, Fri59j]. Weakening [Fri30a]. weapons [Arn03, Fri54b, Fri54i, Fri54c]. Week [Kae39]. Weiss [Fri74a]. Weisskopf [Fri73b]. Wellen [EFS31, EFS32]. Wellenlängen [Fri30a]. went [Ols63]. werden [CSW97]. were [Bey49]. Wheeler [Fri69]. where [Ols63]. Who [MP99, Fri68b, Pei79]. Wild [Pei86]. Windows [AFL39]. Wirkung [Fri27]. Wolf [KED^{+}60]. Woran [Fri81]. Work [Fri79a, Lau51, Mar54, Coc46, Fre86, Roz67, Fri67a]. workers [RRK^{+}64]. Working [Fri65d]. Works [Fri74a, BR64, Fri54d, Fri54e, RRK^{+}64]. World [Stu18, WM15, Ber11, Fri60c, McC01, Sim12, Fri62b]. Wrong [Mei62].

X [Kae39, Bar05, Fri57b, Sta08]. X-Rays [Kae39]. xv [Bar05, Fri73a, Fri73b]. xvi [Fri61d]. xvii [Bar05, FL54]. xxii [Fri74a, Sta08]. xxxiv [Sta08].

Year [FR13, Coc46]. Years [Bar05, Mla98, Kae48, Rot00, Sim14, Sza92a, Sza92b]. York [Bar05, Lin60, Sta08]. young [Fri56e]. Yuan [Fri64e].

Zeit [Fri81]. zum [Fri28]. zur [FPLR59a, KED^{+}60, Fri30b].

References

Arnold:1939:TMW

[AFL39] W. Arnold, O. R. Frisch, and H. Levi. Thin mica windows. Review of scientific instruments, 10(6):197, June 1939. CODEN RSINAK. ISSN 0034-6748 (print), 1089-7623 (electronic).

Alton:1982:RPC

[ALJ82] Jeannine Alton and Julia Latham-Jackson. Report on the papers and correspondence of Otto Robert Frisch, FRS, physicist (1904-1979) deposited in Trinity College Library, Cambridge. Reproduced for the Contemporary Scientific Archives Centre by the Royal Commission on Historical Manuscripts, London, UK, 1982. 155 leaves pp. LCCN Z6611.P57 A47 1982 QC21.2.

Anderson:1973:TQA

[And73] Herbert Anderson. Three questions about the sustained nuclear chain reaction. The University of Chicago Magazine, 65(5):3–7, March/April 1973. ISSN 0041-9508. URL http://library.ucsd.

edu/dc/object/bb3960604v.

Andersen:1998:CSR

[And98] Hanne Andersen. Characteristics of scientific revolutions. Endeavour, 22(1):3–6, January 1998. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016093279801093X.

Anonymous:1939:VEF

[Ano39] Anonymous. Vast energy freed by uranium atom: Split, it produces 2 ‘cannonballs’, each of 100,000,000 electron volts: Hailed as epoch making; new process, announced at Columbia, uses only 1–30 volt to liberate big force. New York Times, ??(??): 18, January 31, 1939. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/102759255/.

Anonymous:1964:BRL

[Ano64] Anonymous. Books received: [… Progress in Nuclear Physics, Volume 9, edited by O. R. Frisch …]. Physics Today, 17(8):54– 57, August 1964. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1973:DNP

[Ano73] Anonymous. Distinguished nuclear pioneer lecturer–1973. Otto Frisch. Journal of Nuclear Medicine, 14(6):372–374, June 1, 1973. CODEN JNMEAQ. ISSN 0161-5505 (print), 1535-5667

(electronic). URL http://jnm.snmjournals.org/content/14/

6/372.full.pdf+html. Reprinted from Trinity Review, Lent Term, 1959, pages 25–27.

Anonymous:1994:EOL

[Ano94] Anonymous. Errata: The Origin of the Liquid-Drop Model and the Interpretation of Nuclear Fission. Perspectives on Science, 2 (2):254, Summer 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Correction of printers error in bottom three equations from page 90.

Aitchison:1977:RPT

[AP77] Ian Johnston Rhind Aitchison and J. E. Paton, editors. Rudolf Peierls and theoretical physics: proceedings of the symposium held in Oxford on July 11th and 12th, 1974, to mark the occasion of the retirement of Professor Sir Rudolph E. Peierls, F.R.S., C.B.E., volume 13 of Progress in nuclear physics. Pergamon Press, New York, NY, USA, 1977. ISBN 0-08-020606-9. LCCN QC1 .P3 1974.

Arms:1966:DP

[Arm66a] Nancy Arms. The diffusion project. In A Prophet in Two Countries: the Life of F. E. Simon [Arm66b], chapter 10, pages 100– 111. ISBN 0-08-011562-4, 0-08-011563-2, 1-4831-6445-4, 1-4831-3915-8 (e-book). LCCN QC16.S5 A7 1966.

Arms:1966:PTC

[Arm66b] Nancy Arms. A Prophet in Two Countries: the Life of F. E. Simon. Pergamon Press, New York, NY, USA, 1966. ISBN 0-08-011562-4, 0-08-011563-2, 1-4831-6445-4, 1-4831-3915-8 (e-book). viii + 171 pp. LCCN QC16.S5 A7 1966.

Arnold:2000:RBK

[Arn00] Lorna Arnold. Recalling Britain’s key nuclear role. Physics World, 13(2):17–18, February 2000. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://stacks.iop.org/

2058-7058/13/i=2/a=18.

Arnold:2003:HNW

[Arn03] Lorna Arnold. The history of nuclear weapons: The Frisch– Peierls memorandum on the possible construction of atomic bombs of February 1940. Cold War History, 3(3):111–126, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682740312331391658.

Baldwin:2014:KRP

[Bal14] Melinda Baldwin. ‘Keeping in the race’: physics, publication speed and national publishing strategies in Nature, 1895–1939. British Journal for the History of Science, 47(2):257–279, June 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://journals.cambridge.org/abstract_

S0007087413000381. Published online, but not yet assigned to a journal volume.

Barschall:1987:RED

[Bar87] Henry H. Barschall. Reminiscences of the early days of fission. Physics Today, 40(6):??, June 1987. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Bartholomew:2005:ERO

[Bar05] James R. Bartholomew. Essay review: One hundred years of the Nobel Science Prizes: Elisabeth Crawford (Editor). Historical Studies in the Nobel Archives: The Prizes in Science and Medicine. viii + 161 pp., Tokyo: Universal Academy Press, 2002. Elisabeth Crawford. The Nobel Population, 1901–1950: A Census of the Nominators and Nominees for the Prizes in Physics and Chemistry. vi + 420 pp., Tokyo: Universal Academy Press, 2002. Mauro Dardo. Nobel Laureates and Twentieth-Century Physics. x + 515 pp., Cambridge: Cambridge University Press, 2004. Robert Marc Friedman. The Politics of Excellence: Behind the Nobel Prize in Science. xv + 400 pp., notes, index. New York: W. H. Freeman, 2001. István Hargittai. The Road to Stockholm: Nobel Prizes, Science, and Scientists. xvii + 342 pp., Oxford/New York: Oxford University Press, 2002. George Thomas Kurian. The Nobel Scientists: A Biographical Encyclopedia. 675 pp., Amherst, N.Y.: Prometheus Books, 2002. Isis, 96(4):625–632, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/498605;
 http://www.jstor.org/stable/3652242.

Bastin:1971:QTB

[Bas71] Ted Bastin, editor. Quantum Theory and Beyond: Essays and Discussions Arising from a Colloquium. Cambridge University Press, Cambridge, UK, 1971. ISBN 0-521-07956-X.

LCCN QC174.1 .Q37. URL http://www.loc.gov/catdir/
enhancements/fy0907/77127237-d.html;   http://www.loc.
gov/catdir/enhancements/fy0907/77127237-t.html.

Bastin:2009:QTB

[Bas09] Ted Bastin, editor. Quantum Theory and Beyond: Essays and Discussions Arising from a Colloquium. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-11548-5, 0-521-07956-X. LCCN QC174.1 .Q83 2009.

Blackett:1946:MAE

[BBD^{+}46] P. M. S. Blackett, M. Born, P. I. Dee, P. A. M. Dirac, N. Feather, E. A. Guggenheim, H. S. W. Massey, P. B. Moon, N. F. Mott, M. L. E. Oliphant, F. A. Paneth, R. E. Peierls, M. H. L. Pryce, F. E. Simon, Sir George Thompson, O. R. Frisch, and H. W. B. Skinner. Memo to the UN Atomic Energy Commission. Bulletin of the Atomic Scientists, 1(12):6–8, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bernstein:2002:HCM

[Ber02] Jeremy Bernstein. Heisenberg and the critical mass. American Journal of Physics, 70(9):911–916, September 2002. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/?AJP/70/911/1.

Bernstein:2011:MCW

[Ber11] Jeremy Bernstein. A memorandum that changed the world. American Journal of Physics, 79(5):440–446, May 2011. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Bethe:1933:TFG

[Bet33] Hans A. Bethe. Theorie des Ferromagnetismus. (German) [Theory of ferromagnetism]. In P. J. W. Debye, editor, Magnetismus, pages 74–81. Hirzel, Leipzig, Germany, 1933.

Beyer:1949:FNP

[Bey49] Robert T. (Robert Thomas) Beyer, editor. Foundations of nuclear physics: facsimiles of thirteen fundamental studies as they were originally reported in the scientific journals. Dover, New York, NY, USA, 1949. 272 pp. LCCN QC173 .B485. See original paper [HS39] and later translation [Gra64].

Bludman:1961:BBN

[BGR61] S. A. Bludman, A. Garren, and A. H. Rosenfeld. Books: The Nation’s Safety and Arms Control, by Arthur T. Hadley; Arms Reduction: Program and Issues, edited by David H. Frisch; Strategy and Arms Control, by Thomas C. Schelling and Morton H. Halperin. Bulletin of the Atomic Scientists, 17(9):380–381, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bhabha:1939:FLI

[Bha39] H. J. Bhabha. The fundamental length introduced by the theory of the mesotron (meson). Nature, 143(3616):276–277, February 18, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/

v143/n3616/pdf/143276b0.pdf. This paper immediately follows [Fri39c], and proposes replacement of the name ‘mesotron’ used by Fermi [FTW47] with the shorter ‘meson’, the form that is now conventional.

Bohr:1939:DHN

[Boh39] Niels Bohr. Disintegration of heavy nuclei. Nature, 143 (3617):330, February 25, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://

www.chemteam.info/Chem-History/Bohr-Fission-1939.html;
 http://www.nature.com/nature/journal/v143/n3617/pdf/
143330a0.pdf.

Bohr:1961:ATD

[Boh61] Niels Bohr. Atomic Theory and the Description of Nature: Four essays, with an introductory survey. Cambridge University Press, Cambridge, UK, 1961. vii + 119 pp. LCCN QC173.18 .B64 1961.

Bohr:1964:NBH

[BR64] Niels Bohr and S. (Stefan) Rozental, editors. Niels Bohr: hans liv og virke fortalt af en kreds af venner og medarbejdere. (Danish) [Niels Bohr: his life and works told by a circle of friends and colleagues]. J. H. Schultz, København, Danmark, 1964. 341 pp. LCCN QC16.B63 N5. English translation in [Roz67].

Brown:1947:BEA

[Bro47] H. Brown. Books: Explaining atoms to the people: The Atomic Story, by John W. Campbell; Meet the Atoms, by O. R. Frisch; Explaining the Atom by Selig Hecht. Bulletin of the Atomic Scientists, 3(6):163, 166, June 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Butler:1964:BRP

[But64] C. C. Butler. Book review: Progress in Nuclear Physics, volume 9, by O. R. Frisch. Oxford: Pergamon Press Ltd. 1963. Pp. v + 310. Physics Bulletin, 15(6):160–161, June 1964. CODEN PHSBB4. ISSN 0031-9112.

Bethe:1980:ORF

[BW80] H. A. Bethe and George Winter. Otto Robert Frisch. Physics Today, 33(1):99–100, January 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.2913924. See correction

[Dit80].

Byers:2002:FS

[Bye02] N. Byers. Fermi and Szilard. ArXiv Physics e-prints, July 2002.

URL http://adsabs.harvard.edu/abs/2002physics...7094B.

Clark:1965:T

[Cla65] Ronald William Clark. Tizard. MIT Press, Cambridge, MA, USA, 1965. xvii + 458 + 8 pp. LCCN Q143.T5 C6. Forewords by Sir Solly Zuckerman and Vannevar Bush.

Cockcroft:1946:RLW

[Coc46] J. D. Cockcroft. Rutherford: Life and work after the year 1919, with personal reminiscences of the Cambridge period. Proceedings of the Physical Society, London, 58(6):625–633, November 1946. CODEN PPSOAU. ISSN 0959-5309 (print), 2051-2171

(electronic). URL http://stacks.iop.org/0959-5309/58/i=6/
a=301.

Crawford:1996:NTW

[CSW96] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. A Nobel tale of wartime injustice. Nature, 382(6590):393–395, August 1, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v382/
n6590/abs/382393a0.html.

Crawford:1997:KIP

[CSW97] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. Die Kernspaltung und ihr Preis. Warum nur Otto Hahn den Nobelpreis erhielt, Otto Frisch, Lise Meitner und Fritz Straßmann dagegen nicht berücksichtigt werden. (German) [Fission and its price. Why only Otto Hahn received the Nobel Prize: Otto Frisch, Lise Meitner and Fritz Strassmann are not taken into consideration]. Kultur & Technik. Zeitschrift des Deutschen Museums München, 21(2):30–35, ???? 1997. ISSN 0344-5690. URL http:/

/www.deutsches-museum.de/fileadmin/Content/data/Insel/
Information/KT/heftarchiv/1997/21-2-30.pdf. Translation

from English by Dieter Beisel of [CSW96].

Deichmann:1996:BUH

[Dei96] Ute Deichmann. Biologists under Hitler. Harvard University Press, Cambridge, MA, USA, 1996. ISBN 0-674-07404-1. xviii + 468 pp. LCCN QH305.2.G3 D4513 1996.

Davies:1970:SFS

[DFS70] D. J. M. Davies, O. R. Frisch, and G. S. B. Street. Sweepnik: A fast semi-automatic track-measuring machine. Nuclear Instruments and Methods, 82:54–60, 1970. CODEN NUIMAL. ISSN 0029-554x (print), 1878-3759 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0029554X70903253.

Ditchburn:1980:FO

[Dit80] R. W. Ditchburn. Frisch obituary. Physics Today, 33(5):86, May 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). Correction to [BW80], noting that Frisch did not succeed to Rutherford’s chair, but rather, was Jacksonian Professor of Natural Philosophy.

Dalitz:1997:SSP

[DP97] Richard Henry Dalitz and Sir Rudolf Peierls, editors. Selected scientific papers of Sir Rudolf Peierls: with commentary, volume 19 of World Scientific series in 20th century physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). xxiii + 805 pp. LCCN QC21.2 .P42 1997. URL

https://ui.adsabs.harvard.edu/#abs/1997sspr.book.....
D;   https://ui.adsabs.harvard.edu/#abs/1997ssps.book.
....P;   https://ui.adsabs.harvard.edu/#abs/1997WSSP..
.19.....D;   https://www.worldscientific.com/worldscibooks/
10.1142/3128.

Duncan:1977:EI

[DWS77] Ronald Duncan and Miranda Weston-Smith, editors. The Encyclopaedia of Ignorance. Pergamon Press, New York, NY, USA, 1977. ISBN 0-08-021238-7 (hardcover), 0-08-022426-1 (flexicover), 0-08-021230-1 (paperback: volume 1), 0-08-021231-X (paperback: volume 2). x + 443 pp. LCCN Q158.5 .E53 1977.

Duncan:1979:LTC

[DWS79] Ronald Duncan and Miranda Weston-Smith, editors. Lying truths: a critical scrutiny of current beliefs and conventions. Pergamon Press, New York, NY, USA, 1979. ISBN 0-08-021978-0. vi + 242 pp. LCCN AZ999 .L94 1979.

Estermann:1931:VMB

[EFS31] I. Estermann, R. Frisch, and O. Stern. Versuche mit monochromatischen de Broglie-Wellen der Molekularstrahlen. (German) [Experiments with monochromatic de Broglie waves of molecular beams]. Zeitschrift für Physik, 72(??):370–674, 1931. CODEN ZEPYAA. ISSN 0044-3328.

Estermann:1932:MBW

[EFS32] I. Estermann, R. Frisch, and O. Stern. Monochromasierung der de Broglie-Wellen von Molekularstrahlen. (German) [Monochromatization of de Broglie waves of molecular beams]. Zeitschrift für Physik, 73(5–6):348–365, May 1932. CODEN ZEPYAA. ISSN 0044-3328.

Estermann:1933:MMP

[EFS33] I. Estermann, R. Frisch, and O. Stern. Magnetic moment of the proton. Nature, 132(3326):169–170, July 29, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v132/
n3326/pdf/132169a0.pdf.

Ellmer:2008:DEK

[Ell08] Reinhold Ellmer. Die drei Entdecker der Kernspaltung. (German) [The three discovers of nuclear fission]. Nachrichten aus der Chemie, 56(12):1241–1243, December 2008. CODEN NACHFB. ISSN 1439-9598 (print), 1868-0054 (electronic). See response [Trö09].

Farmelo:2013:CBHa

[Far13] Graham Farmelo. Churchill’s bomb: a hidden history of science, war and politics. Faber and Faber, London, UK, 2013. ISBN 0-571-24978-7 (hardcover). x + 554 + 12 pp. LCCN ????

Frisch:1960:NEN

[FCJ60] Otto Robert Frisch, Sir J. Cockcroft, and I. G. John. Nuclear energy: its nature, control and use. UNA Peacefinder pamphlet, 29(??):??, ???? 1960.

Feisch:1935:IRF

[Fei35] O. R. Feisch. Induced radioactivity of fluorine and calcium. Nature, 136(3432):220, August 10, 1935. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.

nature.com/articles/136220a0.

Feld:1979:LKO

[Fel79] Bernard T. Feld. Lew Kowarski, 1907–1979. Otto Robert Frisch, 1904–1979. Bulletin of the Atomic Scientists, 35(10):4–5, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feshbach:1954:BRP

[Fes54] Herman Feshbach. Book review: Progress in Nuclear Physics, by Otto R. Frisch. Science, 120(3113):341, August 27, 1954. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://www.jstor.org/stable/10.2307/1683014.

Frisch:1946:NUD

[FF46] Otto Robert Frisch and U. R. Friend. Note on a use of delay lines in counter pulse amplifiers. Report MDDC 238 and LADC 221, Atomic Energy Commission, Oak Ridge, TN, USA, July 6, 1946. 4

pp. URL https://hdl.handle.net/2027/mdp.39015077319377.

Manhattan District declassified code 22 August 1946.

Frisch:1963:EP

[FF63] Otto Robert Frisch and Wilhelm Fucks. Die Elementarteilchen der Physik. (German) [The elementary particles of physics]. Westdeutscher Verlag, Köln, West Germany, 1963. 114 pp. LCCN Q49.C95 A8 Heft 124.

Frisch:1935:PM

[FFPE35] O. R. Frisch, G. Fanselau, A. Prey, and J. Eggert. Physikalische Mitteilungen. (German) [Physical messages]. Die Naturwissenschaften, 23(10):166–168, March 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Frank:1951:LMK

[FFS51] S. G. F. Frank, O. R. Frisch, and G. G. Scarrott. LXIII. A mechanical kick-sorter (pulse size analyser). The London, Edinburgh, and Dublin Philosophical Magazine and Journal of Science, 42(329): 603–611, June 1951.

Frisch:1936:SAN

[FHM36] O. R. Frisch, G. Hevesy, and H. A. C. Mckay. Selective absorption of neutrons by gold. Nature, 137(3456):149–150, January 25, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v137/
n3456/pdf/137149b0.pdf.

Frisch:1931:MVG

[FHS31] R. Frisch, W. Holzer, and Hermann Schröder. Mitteilungen aus verschiedenen Gebieten. (German) [Messages from different areas].

Die Naturwissenschaften, 19(49):991–992, December 1931. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Firk:1964:BRP

[Fir64] F. W. K. Firk. Book review: Progress in Nuclear Physics, Volume 9, edited by O. R. Frisch. Physics Today, 17(11):69–70, November 1964. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Frisch:1954:XPM

[FL54] O. R. Frisch and D. J. Littler. XVII. Pile modulation and statistical fluctuations in piles. The London, Edinburgh, and Dublin Philosophical Magazine and Journal of Science, 45(361):126–140,

February 1954. URL https://www.tandfonline.com/doi/abs/
10.1080/14786440208520432.

Feynman:2006:CFA

[FL06] Richard Phillips Feynman and Ralph Leighton. Classic Feynman: all the adventures of a curious character. W. W. Norton & Co., New York, NY, USA, 2006. ISBN 0-393-06132-9. x + 511 pp.

LCCN QC16.F49 A3 2006. URL ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;   http://www.loc.gov/catdir/
toc/ecip0515/2005018928.html.

Frisch:1928:RPS

[FM28a] Otto Robert Frisch and C. Müller. Registrierendes Prazisionsgerat fur sehr schwache Strome (Lichtintensitaten, Ionisationsvorgange usw.). (German) [Registering precision device for very weak currents (light intensities, ionization processes, etc.)]. Zeitschrift für technische Physik, 9(??):445–451, ???? 1928. CODEN ZTPHAU. ISSN 0373-0093.

Frisch:1928:RWR

[FM28b] Otto Robert Frisch and C. Müller. Über die Realisierung der Warburgschen Rationellen Lichteinheit. (German) [On the realization of the Warburg Rational Light Unit]. In Proc. Int. Congr. on Illumination, pages 1125–?? ????, ????, 1928.

Frisch:1979:O

[FM79] O. R. Frisch and Nevill Mott. Obituary. Contemporary Physics, 20 (6):583, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1959:DCB

[FO59] O. R. Frisch and D. N. Olson. Detection of coherent Bremsstrahlung from crystals. Physical Review Letters, 3(3):141– 142, August 1959. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL https://journals.aps.

org/prl/abstract/10.1103/PhysRevLett.3.141.

Frisch:1960:SAA

[FO60] O. R. Frisch and A. J. Oxley. A semi-automatic analyser for bubble chamber photographs. Nuclear Instruments and Methods, 9(1):92– 96, October 1960. CODEN NUIMAL. ISSN 0029-554X (print), 1878-3759 (electronic).

Ford:2017:BRC

[For17] Peter Ford. Book review: Crystal Clear — The Autobiographies of Sir Lawrence and Lady Bragg. Edited by A. M. Glazer & Patience Thomson OUP 2015. ISBN-13: 978-0-19-874430-6, 448pp, £3. IOP History of Physics Newsletter, 35:54–69, November 2017.

ISSN 1756-168X. URL http://www.iop.org/activity/groups/
subject/hp/newsletter/file_71196.pdf.

Frisch:1931:IIH

[FP31] R. Frisch and Peter Pringsheim. Über die Intensitätsverteilung im Hg-Triplett 2^{3}S_{1}–2^{3}P_{0,1,2} und die mittlere Leuchtdauer der Triplettkomponenten. (German) [On the intensity distribution in the Hg triplet 2^{3}S_{1}–2^{3}P_{0,1,2} and the average life of the triplet components]. Zeitschrift für Physik, 67(3–4):169–178, March 1931. CODEN ZEPYAA. ISSN 0044-3328. URL https://link.

springer.com/article/10.1007/BF01394596.

Frisch:1936:CSN

[FP36] O. R. Frisch and G. Placzek. Capture of slow neutrons. Nature, 137(3461):357, February 29, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v137/n3461/pdf/137357a0.pdf.

Frisch:1939:RSA

[FP39] O. R. Frisch and R. Peierls. Radioactivity and sub-atomic phenomena. Annual Reports on the Progress of Chemistry, 36:7– 32, 1939. CODEN ARPCAW. ISSN 0365-6217 (print), 1754-

7512 (electronic). URL http://pubs.rsc.org/en/content/
articlelanding/1939/ar/ar9393600007.

Frisch:1940:FPM

[FP40a] Otto R. Frisch and Rudolf E. Peierls. The Frisch–Peierls memorandum, part 1. ALSOS Web site., March 1940. URL http://

alsos.wlu.edu/information.aspx?id=3171.

Frisch:1940:MPR

[FP40b] Otto Robert Frisch and Rudolf Peierls. Memorandum on the properties of a radioactive “Super-bomb”. Technical report, University of Birmingham, Birmingham, UK, March

1940. URL http://en.wikipedia.org/wiki/Frisch%E2%80%
93Peierls_memorandum;   http://www.atomicarchive.com/
Bios/Frisch.shtml;   http://www.atomicarchive.com/Bios/
Peierls.shtml;   http://www.atomicarchive.com/Docs/Begin/
FrischPeierls.shtml;   http://www.atomicarchive.com/Docs/
Begin/FrischPeierls2.shtml.

Frisch:1940:CSB

[FP40c] Otto Robert Frisch and Rudolf Peierls. On the construction of a ‘super-bomb’ based on a nuclear chain reaction in uranium. Technical report, University of Birmingham, Birmingham, UK, March 1940. URL http://en.wikipedia.

org/wiki/Frisch%E2%80%93Peierls_memorandum;   http://
www.atomicarchive.com/Bios/Frisch.shtml;   http://www.
atomicarchive.com/Bios/Peierls.shtml;   http://www.atomicarchive.
com/Docs/Begin/FrischPeierls.shtml. Reprinted in [FP64].

Frisch:1940:PRS

[FP40d] Otto Robert Frisch and Rudolf Peierls. The properties of a radioactive “Super-bomb”. Memorandum, University of Birmingham, Birmingham, UK, March 1940. URL http://en.wikipedia.

org/wiki/Frisch%E2%80%93Peierls_memorandum;   http://
www.atomicarchive.com/Bios/Frisch.shtml;   http://www.
atomicarchive.com/Bios/Peierls.shtml;   http://www.atomicarchive.
com/Docs/Begin/FrischPeierls.shtml;   http://www.atomicarchive.
com/Docs/Begin/FrischPeierls2.shtml. Reprinted in [Cla65].

Frisch:1943:LAT

[FP43] Otto Robert Frisch and Rudolf Peierls. Light absorption by two black hemispheres. Report MS-77 (PRO AB4/915), Directorate of Tube Alloys, ????, UK, 1943.

Frisch:1964:CSB

[FP64] Otto Robert Frisch and Rudolf Peierls. On the construction of a ‘super-bomb’ based on a nuclear chain reaction in uranium [Part 1]. In Britain and Atomic Energy, 1939–1945 [Gow64], pages 389– 393. LCCN QC773.A1 G6 1964. With an introductory chapter by Kenneth Jay.

Frisch:1965:CSB

[FP65] Otto Robert Frisch and Rudolf Peierls. On the construction of a ‘super-bomb’ based on a nuclear chain reaction in uranium [Part 2]. In Tizard [Cla65], pages 215–217. LCCN Q143.T5 C6. Forewords by Sir Solly Zuckerman and Vannevar Bush.

Frisch:1997:FPM

[FP97] O. R. Frisch and R. Peierls. The Frisch–Peierls memorandum of 1940. In Dalitz and Peierls [DP97], pages 277–282. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.adsabs.

harvard.edu/#abs/1997sspr.book..277F.

Frisch:2004:MPR

[FP04] Otto R. Frisch and Rudolf Peierls. Memorandum on the properties of a radioactive super-bomb. In Kelly [Kel04], pages 177–180. ISBN 981-256-040-8. LCCN QC773.A1 R46 2004.

Frisch:1959:BPC

[FPLR59a] O. R. Frisch, F. A. Paneth, F. Laves, and P. Rosbaud, editors. Beiträge zur Physik und Chemie des 20. Jahrhunderts. (German) [Contributions to the physics and chemistry of the 20th century]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1959. 285 pp. LCCN QC475 .B45 1959.

Frisch:1959:TAP

[FPLR59b] O. R. Frisch, F. A. Paneth, F. Laves, and P. Rosbaud, editors. Trends in atomic physics; essays dedicated to Lise Meitner, Otto Hahn, Max von Laue on the occasion of their 80th birthday: Atomic physics. Interscience Publishers, New York, NY, USA, 1959. 285 pp. LCCN QC475 .B45 1959a. An identical edition is published simultaneously under the title Beiträge zur Physik und Chemie des 20. Jahrhunderts.

Frisch:1932:PSQ

[FPSS32] R. Frisch, T. E. Phipps, E. Segrè, and O. Stern. Process of space quantisation. Nature, 130(3293):892–893, December 10, 1932. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v130/
n3293/pdf/130892d0.pdf.

Fernandez:2013:UMA

[FR13] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (ebook). xviii + 522 pp. LCCN QC773 .F47 2013.

French:1986:NBH

[Fre86] A. P. French. Niels Bohr at 100: his life and work. Physics Education, 21(4):220–226, July 1986. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http://stacks.iop.

org/0031-9120/21/i=4/a=005.

Frisch:1926:VSD

[Fri26] Otto Robert Frisch. Verfärbung von Steinsalz durch Kathodenstrahlen. (German) [Discoloration of rock salt by cathode rays]. Dr.Phil. thesis, Institut für Radiumforschung, Universität Wien, Vienna, Austria, 1926.

Frisch:1927:WLK

[Fri27] Otto Robert Frisch. Wirkung von langsamen Kathodenstrahlen auf Steinsalz. (German) [Effect of slow cathode rays on rock salt]. Sber. Akad. Wiss. Wien, 136(??):57–64, ???? 1927.

Frisch:1928:GAS

[Fri28] R. Frisch. Ein Gerät zum Ausmessen von Spektralphotographien, Registrieraufnahmen und dergleichen. (German) [A device for measuring spectral photographs, recording images and the like]. Zeitschrift für Physik, 49(7–8):608, July 1928. CODEN ZEPYAA.

ISSN 0044-3328. URL https://link.springer.com/article/
10.1007/BF01333645.

Frisch:1930:SSV

[Fri30a] Otto Robert Frisch. Schwachung von Strahlen verschiedener Wellenlängen durch Tau-Schichten. (German) [Weakening of rays of different wavelengths by tau layers]. Unknown, 1930.

Frisch:1930:DLG

[Fri30b] R. Frisch. Zur Drehimpulsbilanz bei Lichtemissionsvorgängen. (German) [Toward the angular momentum balance in light emission processes]. Zeitschrift für Physik, 61(9–10):626–631, September 1930. CODEN ZEPYAA. ISSN 0044-3328.

Frisch:1932:EIP

[Fri32] R. Frisch. Elektronenbeugung und inneres Potential der Metalle. (German) [Electron diffraction and internal potential of metals]. Die Naturwissenschaften, 20(37):689, September 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Frisch:1933:ARB

[Fri33a] R. Frisch. Anomalien bei der Reflexion und Beugung von Molekularstrahlen an Kristallspaltflächen. II. (German) [Anomalies in specular reflection and diffraction of molecular beams at crystal slits. II]. Zeitschrift für Physik, 84(7–8):443–447, July 1933. CO-DEN ZEPYAA. ISSN 0044-3328. URL https://link.springer.

com/article/10.1007/BF01342224.

Frisch:1933:ENE

[Fri33b] R. Frisch. Experimenteller Nachweis des Einsteinschen Strahlungsrückstoßes. (German) [Experimental proof of Einstein’s radiation recoil]. Zeitschrift für Physik, 86(1–2):42–48, January 1933. CODEN

ZEPYAA. ISSN 0044-3328. URL https://link.springer.com/
article/10.1007/BF01340182.

Frisch:1934:IRS

[Fri34] O. R. Frisch. Induced radioactivity of sodium and phosphorus. Nature, 133(3367), May 12, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v133/n3367/pdf/133721b0.pdf.

Frisch:1937:SCS

[Fri37] Otto Robert Frisch. On the selective capture of slow neutrons. Det Kongelige Danske Videnskabernes Selskab. Mathematisk-fysiske Meddelelser, 14(12):1–31, ???? 1937.

Frisch:1939:SCC

[Fri39a] O. R. Frisch. Statistical calculation of composite decay curves. Nature, 143(3629):852–853, May 20, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v143/n3629/pdf/143852b0.pdf.

Frisch:1939:RSP

[Fri39b] Otto R. Frisch. Radioactivity and subatomic phenomena. Annual Reports on the Progress of Chemistry, 36:7–24, 1939. CODEN ARPCAW. ISSN 0365-6217 (print), 1754-7512 (electronic).

Frisch:1939:PED

[Fri39c] Otto Robert Frisch. Physical evidence for the division of heavy nuclei under neutron bombardment. Nature, 143(3616):276, February 18, 1939. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.chemteam.info/Chem-
History/Frisch-Fission-1939.html;   http://www.nature.
com/nature/journal/v143/n3616/pdf/143276a0.pdf.

Frisch:1940:RSP

[Fri40] O. R. Frisch. Radioactivity and subatomic phenomena. Annual Reports on the Progress of Chemistry, 37:7–22, 1940. CODEN ARPCAW. ISSN 0365-6217 (print), 1754-7512 (electronic).

Frisch:1941:RSA

[Fri41] O. R. Frisch. Radioactivity and sub-atomic phenomena. Annual Reports on the Progress of Chemistry, 38:287–297, 1941. CODEN ARPCAW. ISSN 0365-6217 (print), 1754-7512 (electronic).

Frisch:1942:ACN

[Fri42a] Otto Robert Frisch. Absolute calibration of a neutron source. Liverpool Report AB 4/50 (BR 49), Directorate of Tube Alloys, ????, UK, 1942.

Frisch:1942:IAU

[Fri42b] Otto Robert Frisch. Isotope analysis of uranium samples by means of α-ray groups. Liverpool Report AB 4/50 (BR 49), Directorate of Tube Alloys, ????, UK, 1942. URL http://discovery.

nationalarchives.gov.uk/details/r/C1978077.

Frisch:1943:RSA

[Fri43] O. R. Frisch. Radioactivity and sub-atomic phenomena. Annual Reports on the Progress of Chemistry, 40:5–11, 1943. CODEN ARPCAW. ISSN 0365-6217 (print), 1754-7512 (electronic).

Frisch:1946:TNP

[Fri46] Otto Robert Frisch. The tools of nuclear physics. Penguin Science News, 2(??):128–161, ???? 1946.

Frisch:1947:LNP

[Fri47a] Otto Robert Frisch. Lectures on nuclear physics at A.E.R.E. Unknown., 1947.

Frisch:1947:MAPa

[Fri47b] Otto Robert Frisch. Meet the atoms; a popular guide to modern physics. Sigma Books, London, UK, 1947. xiv + 226 pp. LCCN QC173 .F723 1947a.

Frisch:1947:MAPb

[Fri47c] Otto Robert Frisch. Meet the atoms, a popular guide to modern physics. A. A. Wyn, Inc., New York, NY, USA, 1947. xiv + 226 pp. LCCN QC173 .F723.

Frisch:1948:MTR

[Fri48a] Otto Robert Frisch. Meson theory and radar. Technion Yb., 1948.

Frisch:1948:SMP

[Fri48b] Otto Robert Frisch. Statistics of multiplicative processes. Report ????, Atomic Energy Research Establishment, Harwell, Berkshire, UK, ???? 1948. ?? pp.

Frisch:1950:BRA

[Fri50a] Otto Robert Frisch. Book review: Acceleration of particles to high energies, London: Institute of Physics. Discovery, 11(??):340, ???? 1950. CODEN DISCAH. ISSN 0012-3625.

Frisch:1950:CL

[Fri50b] Otto Robert Frisch. The cyclotron and its limitations. Discovery, 11(??):262–266, ???? 1950. CODEN DISCAH. ISSN 0012-3625.

Frisch:1950:PNP

[Fri50c] Otto Robert Frisch, editor. Progress in Nuclear Physics. Pergamon Press, New York, NY, USA, 1950. CODEN PNUPAT. ISSN 0079-659X. LCCN QC770 .P76. Frisch was the long-time editor for volumes 2–9 (1953–1960) of this series (volume 1 (1953)–volume 13 (1977)).

Frisch:1950:SC

[Fri50d] Otto Robert Frisch. Scintillation counters. Helvetica Physica Acta, 23(Supplement 3):150–154, 154 1950. CODEN HPACAK. ISSN

0018-0238. URL https://www.e-periodica.ch/digbib/view?
pid=hpa-001:1950:23::1114.

Frisch:1951:AAAb

[Fri51a] O. R. Frisch. Artificial acceleration of atomic particles. Nature, 168(4281):849–851, November 17, 1951. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v168/n4281/pdf/168849a0.pdf.

Frisch:1951:AR

[Fri51b] O. R. Frisch. Atomic reminiscences. Nature, 168(4262):6, July 7, 1951. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v168/
n4262/pdf/168006b0.pdf.

Frisch:1951:AAAa

[Fri51c] Otto Robert Frisch. The artificial acceleration of atomic particles. Proceedings of the Royal Institution of Great Britain, ??(??):??, May 11, 1951. CODEN PIGBAI. ISSN 0035-8959.

Frisch:1951:BRN

[Fri51d] Otto Robert Frisch. Book review: New Atoms, by Otto Hahn. Nature, 168(??):67, ???? 1951. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Frisch:1953:CMP

[Fri53a] Otto Robert Frisch. Causality in modern physics. The Listener, 49(??):138–142, ???? 1953.

Frisch:1953:GAW

[Fri53b] Otto Robert Frisch. Giant accelerators: what do we expect of them? Times Science Review, ??(??):2, Summer 1953.

Frisch:1954:FCD

[Fri54a] O. R. Frisch. On the feasibility of coal-driven power stations. Bulletin of the Atomic Scientists, 10(6):224, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frisch:1954:AWa

[Fri54b] Otto Robert Frisch. Atomic weapons. Atomic Scientists Journal, 8(??):193, ???? 1954.

Frisch:1954:AWc

[Fri54c] Otto Robert Frisch. Atomic weapons. In Rotblat [Rot54], page ?? Published for Atomic Scientist’ Association.

Frisch:1954:HAE

[Fri54d] Otto Robert Frisch. How atomic energy works. The Sunday Times, ??(??):??, October 6, 1954.

Frisch:1954:HHB

[Fri54e] Otto Robert Frisch. How the hydrogen bomb works. The Listener, 51(??):907–908, ???? 1954.

Frisch:1954:IMP

[Fri54f] Otto Robert Frisch. Individuality in modern physics. The Listener, 51(??):57–58, ???? 1954.

Frisch:1954:SHB

[Fri54g] Otto Robert Frisch. Scientists and the hydrogen bomb. The Listener, 51(??):556, ???? 1954.

Frisch:1954:AEH

[Fri54h] Professor O. R. Frisch, O.B.E., F.R.S. Atomic energy — how it all began. British Journal of Applied Physics, 5(3):81–84, March 1954. CODEN BJAPAJ. ISSN 0508-3443 (print), 2057-7656 (elec-

tronic). URL http://iopscience.iop.org/0508-3443/5/3/
301;   http://stacks.iop.org/0508-3443/5/i=3/a=301. A

lecture delivered in London to the Education Group of The Institute of Physics on 20 October, 1953.

Frisch:1954:AWb

[Fri54i] Professor O. R. Frisch, O.B.E., F.R.S. Atomic weapons. Nature, 173(4402):477, March 13, 1954. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v173/n4402/abs/173477a0. From the sec-

ond of a series of public lectures arranged by the Atomic Scientists’ Association and Department of Extra-Mural Studies of the University of London.

Frisch:1955:IRA

[Fri55a] O. R. Frisch. II. — La radioactivité artificielle et la physique. (French) [II. — Artificial radioactivity and physics]. Journal de Physique et le Radium, 16(10):748–753, October 1955. CODEN JPRAAJ. ISSN 0368-3842. Lecture at the celebration of the 25th anniversary of the discovery of artificial radioactivity.

Frisch:1955:PEF

[Fri55b] O. R. Frisch. Prof. Enrico Fermi, For. Mem. R. S. Nature, 175 (4444):18–19, January 1, 1955. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v175/n4444/pdf/175018a0.pdf.

Frisch:1955:AP

[Fri55c] Otto Robert Frisch. The anti-proton. Discovery, 51(??):498, ???? 1955. CODEN DISCAH. ISSN 0012-3625.

Frisch:1955:BR

[Fri55d] Otto Robert Frisch. Book review: Atoms in the Family: My Life with Enrico Fermi — Designer of the First Atomic Pile, by Laura Fermi. Pp. 284 + 15 plates (London: George Allen and Unwin, Ltd., 1955). Nature, 176(4488):850, November 5, 1955. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v176/
n4488/pdf/176850b0.pdf.

Frisch:1955:CAP

[Fri55e] Otto Robert Frisch. Counting atomic particles. Times Science Review, ??(??):6, Spring 1955.

Frisch:1955:HEA

[Fri55f] Otto Robert Frisch. High-energy accelerators. The synchrotron. Discovery, 16(??):450, ???? 1955. CODEN DISCAH. ISSN 0012-3625.

Frisch:1956:JP

[Fri56a]  O. R. Frisch.  “Jocular physics”.  Scientific American,

194(3):93–103, March 1956. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v194/n3/pdf/
scientificamerican0356-93.pdf.

Frisch:1956:APA

[Fri56b] Otto Robert Frisch. Anti-particles and anti-matter. Times Science Review, ??(??):1, Winter 1956.

Frisch:1956:AP

[Fri56c] Otto Robert Frisch. The anti-proton. Nucl. Pwr, 1(??):51–52, ???? 1956.

Frisch:1956:DAP

[Fri56d] Otto Robert Frisch. The discovery of the anti-proton. The Listener, 55(??):16–17, ???? 1956.

Frisch:1956:LYS

[Fri56e] Otto Robert Frisch. Letter to a young scientist. The Listener, 56 (??):269, 350, ???? 1956.

Frisch:1956:FCD

[Fri56f] Otto Robert Frisch. On the feasibility of coal-driven power stations. Nucl. Engng, Lond, ??(??):268–269, ???? 1956.

Frisch:1957:BR

[Fri57a] O. R. Frisch. [Book reviews]. Nature, 179(4564), April 20, 1957. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v179/
n4564/pdf/179799c0.pdf.

Frisch:1957:BRN

[Fri57b] Otto Robert Frisch. Book review: Nuclear explosions and their effects. Pp. x + 184. (Delhi: Indian Ministry of Information & Broadcasting, 1956). Nature, 179(4564):799, April 20, 1957. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Frisch:1957:PEW

[Fri57c] Otto Robert Frisch. Die Probleme der Energieerzeugung aus Wasserstoff. (German) [The Problems of Hydrogen Generation]. Österreichische Studiengesellschaft für Atomenergie Ges. m.b.H. (SGAE), Vienna, Austria, 1957. ??–?? pp.

Frisch:1957:HNF

[Fri57d] Otto Robert Frisch. How nuclear fission came to be discovered. New Scientist, 2(??):29–31, ???? 1957.

Frisch:1957:PC

[Fri57e] Otto Robert Frisch. Parity is not conserved. University Quarterly, 11(??):235–244, ???? 1957.

Frisch:1957:PUA

[Fri57f] Otto Robert Frisch. Peaceful uses of atomic fission. Atlantic Monthly, ??(??):??, ???? 1957. ISSN 1072-7825 (print), 2151-9463 (electronic).

Frisch:1957:QAM

[Fri57g] Otto Robert Frisch. The quest for anti-matter. New Scientist, 1 (??):37–39, January 31, 1957. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Frisch:1958:NH

[Fri58a] Otto Robert Frisch. The nuclear handbook. Van Nostrand, Princeton, NJ, USA, 1958. ???? pp. LCCN QC783 .F7. With 22 specialist contributors.

Frisch:1958:NHS

[Fri58b] Otto Robert Frisch. The nuclear handbook. G. Newnes, London, UK, 1958. ???? pp. LCCN QC783 .F7 1958a. With 22 specialist contributors.

Frisch:1959:MB

[Fri59a] O. R. Frisch. Molecular beams. Contemporary Physics, 1(1):3– 16, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1959:PCN

[Fri59b] O. R. Frisch. “Parity is not conserved” a new twist to physics? Bulletin of the Atomic Scientists, 15(3):139–143, March 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Universities Quarterly.

Frisch:1959:BKB

[Fri59c] Otto Robert Frisch. Betrachtungen über kohärente Bremsstrahlung. (German) [Considerations for coherent Bremsstrahlung]. Acta Physica Austriaca, 12(??):331–335, ???? 1959. CODEN APASAP. ISSN 0001-6713.

Frisch:1959:BRT

[Fri59d] Otto Robert Frisch. Book review: Turning points in physics (Oxford lectures). New Scientist, 6(??):527–528, ???? 1959. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Frisch:1959:BRC

[Fri59e] Otto Robert Frisch. Book review: Čerenkov radiation and its application, by C. V. Jelley. New Scientist, 5(??):1034, ???? 1959. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Frisch:1959:EN

[Fri59f] Otto Robert Frisch. The energetic nucleus. United Nations News, ??(??):7–10, January/March 1959.

Frisch:1959:MA

[Fri59g] Otto Robert Frisch. Measuring atoms. The Listener, 62(??):1161– 1163, ???? 1959.

Frisch:1959:NEP

[Fri59h] Otto Robert Frisch. The new elementary particles. Times Science Review, ??(??):12, Autumn 1959.

Frisch:1959:NP

[Fri59i] Otto Robert Frisch. The new particles. Bulletin of the Institute of Physics, London, 10(??):137–142, ???? 1959.

Frisch:1959:PNC

[Fri59j] Otto Robert Frisch. Parity non-conservation in weak interactions. Progress in Nuclear Physics, 6(??):267–273, ???? 1959. CODEN PNUPAT. ISSN 0079-659X.

Frisch:1959:PAN

[Fri59k] Otto Robert Frisch. Probing the atomic nucleus. New Scientist, 6 (??):228–230, 416, ???? 1959.

Frisch:1959:AM

[Fri59l] Otto Robert Frisch. This age of magnetism. The Listener, 61(??): 1108–1109, ???? 1959.

Frisch:1960:LNSa

[Fri60a] Otto Robert Frisch. Are the laws of nature symmetrical? Proceedings of the Royal Institution of Great Britain, 38(??):249–??, ???? 1960. CODEN PIGBAI. ISSN 0035-8959.

Frisch:1960:BRG

[Fri60b] Otto Robert Frisch. Book review: The great decision, by M. Amrine. New Scientist, 7(??):1555–1556, ???? 1960.

Frisch:1960:ESA

[Fri60c] Otto Robert Frisch. Exploring the sub-atomic world. The Listener, 63(??):119–120, ???? 1960.

Frisch:1960:FP

[Fri60d] Otto Robert Frisch. Fundamental particles. In Haslett and John [HJ60], pages 9–22.

Frisch:1960:SPa

[Fri60e] Otto Robert Frisch. The strange particles. The Listener, 63(??): 173–176, ???? 1960.

Frisch:1960:SPb

[Fri60f] Otto Robert Frisch. Strangeness and parity. The Listener, 63(??): 217–218, ???? 1960.

Frisch:1960:LNSb

[Fri60g] O. R. Frischb. Are the laws of nature symmetrical? Nature, 187 (4738):638–641, August 20, 1960. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v187/n4738/pdf/187638a0.pdf.

Frisch:1961:PNP

[Fri61a] O. R. Frisch. Progress in nuclear physics, volume 3. ????, ????, 1961. vii + 279 pp.

Frisch:1961:TRP

[Fri61b] O. R. Frisch. Time and relativity: Part I. Contemporary Physics, 3(1):16–27, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1961:APT

[Fri61c] Otto Robert Frisch. Atomic physics today. Basic Books, New York, NY, USA, 1961. vi + 254 pp. LCCN QC778 .F73.

Frisch:1961:BRD

[Fri61d] Otto Robert Frisch. Book review: From dualism to unity in physics. By A. Landé (Cambridge University Press, 1960). Pp. xvi + 114. Contemporary Physics, 2(4):323, 1961. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1961:BRB

[Fri61e] Otto Robert Frisch. Book review: The birth of the bomb, Robin Clark. Discovery, 22(??):318, ???? 1961. CODEN DISCAH. ISSN 0012-3625.

Frisch:1961:EP

[Fri61f] Otto Robert Frisch. The elementary particles. Discovery, 22(??): 515–524, ???? 1961. CODEN DISCAH. ISSN 0012-3625.

Frisch:1961:I

[Fri61g] Otto Robert Frisch. Introduction. In Atomic Theory and the Description of Nature: Four essays, with an introductory survey [Boh61], page ?? LCCN QC173.18 .B64 1961.

Frisch:1961:FP

[Fri61h] Professor O. R. Frisch, O.B.E., D.Sc., F.R.S. Fundamental particles. Journal of the Institution of Electrical Engineers, 7(78): 357–359, June 1961. CODEN JISEAL. ISSN 0368-2692 (print), 2054-0574 (electronic).

Frisch:1962:TRP

[Fri62a] O. R. Frisch. Time and relativity: Part II. Contemporary Physics, 3(3):194–201, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1962:BRN

[Fri62b] Otto Robert Frisch. Book review: The new world, by R. G. Hewlett & O. E. Anderson. New Scientist, 15(??):46–??, ???? 1962. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Frisch:1962:BRP

[Fri62c] Otto Robert Frisch. Book review: The philosophical impact of contemporary physics, M. Čapek. Discovery, 23(??):44, ???? 1962. CODEN DISCAH. ISSN 0012-3625.

Frisch:1962:OQ

[Fri62d] Otto Robert Frisch. Observation and the quantum. ????, ??(??): 309, ???? 1962.

Frisch:1963:MP

[Fri63a] Otto Robert Frisch. The magnetic proton. The Listener, 70(??): 459–460, ???? 1963.

Frisch:1963:MOF

[Fri63b] Otto Robert Frisch. Means of observing fundamental particles. In Garratt [Gar63a], page ??

Frisch:1964:BRI

[Fri64a] O. R. Frisch. Book review: An introduction to the theory of relativity, by A. V. Rosser. Contemporary Physics, 6(2):155–156, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1964:NNR

[Fri64b] O. R. Frisch. Nuclei and nuclear reactions. Contemporary Physics, 6(2):81–93, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1964:TNP

[Fri64c] Otto R. Frisch. The tools of nuclear physics. In Rapport and Wright [RW64], pages 144–176.

Frisch:1964:BRP

[Fri64d] Otto Robert Frisch. Book review: Physics in the sixties, edited by S. K. Runcorn. Discovery, 24(??):50, ???? 1964. CODEN DISCAH. ISSN 0012-3625.

Frisch:1964:BRN

[Fri64e] Otto Robert Frisch. Book review: The nature of matter, by Yuan. ????, ??(??):??, ???? 1964.

Frisch:1964:C

[Fri64f] Otto Robert Frisch. Causality. The Listener, 72(??):83–84, ???? 1964.

Frisch:1964:ISS

[Fri64g] Otto Robert Frisch. Interessen samler sig omkring atomkernen. (Danish) [Interest collects around the atomic nucleus]. In Bohr and Rozental [BR64], pages 132–144. LCCN QC16.B63 N5. English translation in [Roz67].

Frisch:1964:NFP

[Fri64h] Otto Robert Frisch. The new fundamental particles. The Listener, 71(??):119–120, 173, 176, 217–218, ???? 1964.

Frisch:1965:MB

[Fri65a] O. R. Frisch. Molecular beams. Scientific American, 212 (5):58–74, May 1965. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v212/n5/pdf/scientificamerican0565-
58.pdf.

Frisch:1965:TPL

[Fri65b] O. R. Frisch. Take a photon . . .. Contemporary Physics, 7(1):45– 53, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1965:PMM

[Fri65c] Otto Robert Frisch. Poor man’s SMP. IEEE Transactions on Nuclear Science, 12(4):196–197, August 1965. CODEN IRNSAM. ISSN 0018-9499 (print), 1558-1578 (electronic).

Frisch:1965:WA

[Fri65d] Otto Robert Frisch. Working with atoms. Basic Books, New York, NY, USA, 1965. 96 pp.

Frisch:1965:PMP

[Fri65e] Professor O. R. Frisch, O.B.E., F.R.S. The particles of modern physics. Electronics and Power, 11(5):153, May 1965. CODEN ELPWAQ. ISSN 0013-5127 (print), 2053-7883 (electronic).

Frisch:1966:SLD

[Fri66a] Otto Robert Frisch. Symmetry laws in dispute. The Listener, 76 (??):725, ???? 1966.

Frisch:1966:PMP

[Fri66b] Professor O. R. Frisch, O.B.E., F.R.S. The particles of modern physics. Proceedings of the Institution of Electrical Engineers, 113 (3):529–537, March 1966. CODEN PIEEAH. ISSN 0020-3270 (print), 2053-7891 (electronic). The Fifty-Sixth Kelvin Lecture.

Frisch:1967:BNB

[Fri67a] O. R. Frisch. Books: Niels Bohr, his life and work, edited by S. Rozental. Niels Bohr, by Ruth Moore. Scientific American, 216(6):145–153, June 1967. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v216/n6/pdf/
scientificamerican0667-145.pdf.

Frisch:1967:IFA

[Fri67b] O. R. Frisch. The interest is focussing on the atomic nucleus. In Rozental [Roz67], pages 137–148. LCCN QC16.B63 N53.

Frisch:1967:TG

[Fri67c] Otto Robert Frisch. Tracks galore. Camb. Res., 2(??):3, ???? 1967.

Frisch:1967:WP

[Fri67d] Otto Robert Frisch. What is physics? B.B.C. Publications., ?? (??):??, ???? 1967.

Frisch:1968:OH

[Fri68a] O. R. Frisch. Otto Hahn. Physics Bulletin, 19(10):354, October 1968. CODEN PHSBB4. ISSN 0031-9112.

Frisch:1968:SWO

[Fri68b] O. R. Frisch. Scientist who open the way to atom bomb. The Observer, ??(??):??, August 4, 1968.

Frisch:1968:LMD

[Fri68c] Otto R. Frisch. Lise Meitner dies; nuclear-physics pioneer. Physics Today, 21(12):101, December 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

//scitation.aip.org/content/aip/magazine/physicstoday/
article/21/12/10.1063/1.3034630.

Frisch:1969:BRE

[Fri69] Otto Robert Frisch. Book review: Einstein’s vision, by J. A. Wheeler. Camb. Res., 5(3):75, ???? 1969.

Frisch:1970:FNE

[Fri70a] Otto Robert Frisch. The first nuclear explosion. New Scientist, 47(??):274–278, ???? 1970. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Frisch:1970:LM

[Fri70b] Otto Robert Frisch. Lise Meitner. 1878–1968. Biographical Memoirs of Fellows of the Royal Society, 16:405–420, November 1970.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL http://www.jstor.org/stable/769597.

Frisch:1971:CPQ

[Fri71] Otto Robert Frisch. The conceptual problem of quantum theory from the experimentalist’s point of view. In Bastin [Bas71], chapter 2, pages 13–21. ISBN 0-521-07956-X. LCCN QC174.1 .Q37.

URL http://www.loc.gov/catdir/enhancements/fy0907/77127237-
d.html;   http://www.loc.gov/catdir/enhancements/fy0907/
77127237-t.html.

Frisch:1972:MBC

[Fri72a] Otto Robert Frisch. Measuring bubble-chamber tracks. In ????, editor, Conference on Machine Perception, volume 13, page ?? IOP Publishing, Bristol, UK, 1972.

Frisch:1972:NM

[Fri72b] Otto Robert Frisch. The Nature of Matter. The World of science library. Thames and Hudson, London, UK, 1972. ISBN 0-500-08006-2, 0-500-10006-3 (paperback). 216 pp. LCCN QC173 .F724.

Frisch:1973:BRA

[Fri73a] O. R. Frisch. Book review: Albert Einstein: Creator and Rebel, by B. Hoffman. London: Hart-Davis, MacGibbon, 1973. pp xv + 272. Physics Bulletin, 24(11):679, November 1973. CODEN PHSBB4. ISSN 0031-9112.

Frisch:1973:BRP

[Fri73b] O. R. Frisch. Book review: Physics in the 20th Century, by V. F. Weisskopf. London: MIT Press, 1972, pp xv + 368. Physics Bulletin, 24(8):491, August 1973. CODEN PHSBB4. ISSN 0031-9112.

Frisch:1973:MBC

[Fri73c] O. R. Frisch. Measuring bubble chamber tracks. Physics Bulletin, 24(5):275–277, May 1973. CODEN PHSBB4. ISSN 0031-9112.

Frisch:1973:DNP

[Fri73d] Otto Robert Frisch. Distinguished nuclear pioneer–1973. Lise Meitner. Journal of Nuclear Medicine, 14(6):365–371, June 1, 1973. CODEN JNMEAQ. ISSN 0161-5505 (print), 1535-5667

(electronic). URL http://jnm.snmjournals.org/content/14/

6/365.full.pdf+html. Reprinted from The Biographical Memoirs of Fellows of The Royal Society, Volume 16, November, 1970, pages 405–416.

Frisch:1973:NM

[Fri73e] Otto Robert Frisch. The Nature of Matter. The World of science library. Dutton, New York, NY, USA, 1973. 216 pp. LCCN QC173 .F724.

Frisch:1974:BRC

[Fri74a] O. R. Frisch. Book review: The Collected Works of Leo Szilard: Scientific Papers, by Bernard T. Feld and Gertrude Weiss Szilard (eds.), London: MIT Press 1973, pp. xxii + 737. Physics Bulletin, 25(1):27, January 1974. CODEN PHSBB4. ISSN 0031-9112.

Frisch:1974:STS

[Fri74b] Otto R. Frisch. Somebody turned the Sun on with a switch. Bulletin of the Atomic Scientists, 30(4):12–18, April 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frisch:1974:DMT

[Fri74c] Otto Robert Frisch. Distinguishing the main types of radioactivity. [Sound recording]. Spring Green Multimedia, Washington, DC, USA, 1974. LCCN QC795; RYB 6353. 1 cassette tape.

Frisch:1974:LM

[Fri74d] Otto Robert Frisch. Lise Meitner. In Charles Coulston Gillispie, editor, Dictionary of scientific biography, pages 260–263. Scribner, New York, NY, USA, 1974.

Frisch:1974:ONF

[Fri74e] Otto Robert Frisch. The origin of nuclear fission. [Sound recording]. Spring Green Multimedia, Washington, DC, USA, 1974. LCCN QC790; RYB 6354. 1 cassette tape.

Frisch:1974:PP

[Fri74f] Otto Robert Frisch. Presenting physics. Physics Bulletin, 25(9):392, September 1974. CODEN PHSBB4. ISSN 0031-

9112. URL http://iopscience.iop.org/article/10.1088/
0031-9112/25/9/023;   http://stacks.iop.org/0031-9112/
25/i=9/a=023.

Frisch:1975:OSG

[Fri75a] O. R. Frisch. Obituary: Sir George Paget Thomson F.R.S. (1892– 1975). The Philosophical Magazine: A Journal of Theoretical Experimental and Applied Physics, 32(6):0, December 1975.

Frisch:1975:SSD

[Fri75b] Otto Robert Frisch. Shooting sparrows in the dark. New Scientist, 65(??):654–657, ???? 1975.

Frisch:1975:WS

[Fri75c] Otto Robert Frisch. A walk in the snow. New Scientist, 60(??): 833, ???? 1975.

Frisch:1977:BRQ

[Fri77a] Otto Robert Frisch. Book review: From quarks to quasars, by E. Thomas. New Scientist, 74(??):481–482, ???? 1977.

Frisch:1977:EST

[Fri77b] Otto Robert Frisch. Early steps towards the chain reaction. In Aitchison and Paton [AP77], pages 18–27. ISBN 0-08-020606-9. LCCN QC1 .P3 1974.

Frisch:1977:FSA

[Fri77c] Otto Robert Frisch. The first sub-atomic particle. New Scientist, 76(??):408–410, ???? 1977. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Frisch:1977:W

[Fri77d] Otto Robert Frisch. Why? In Duncan and Weston-Smith [DWS77], pages 1–4. ISBN 0-08-021238-7 (hardcover), 0-08-022426-1 (flexicover), 0-08-021230-1 (paperback: volume 1), 0-08-021231-X (paperback: volume 2). LCCN Q158.5 .E53 1977.

Frisch:1978:IT

[Fri78] Otto Robert Frisch. Impossible things. New Scientist, 78(??): 688–689, ???? 1978. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Frisch:1979:EWN

[Fri79a] Otto R. Frisch. Experimental work with nuclei: Hamburg, London, Copenhagen. In Stuewer [Stu79], pages 65–79. ISBN 0-8166-0869-5. LCCN QC773 .S95 1977.

Frisch:1979:PAA

[Fri79b] Otto R. Frisch. Prometheus of the atomic age. Nature, 282(5735), November 8, 1979. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v282/n5735/pdf/282136a0.pdf.

Frisch:1979:GRH

[Fri79c] Otto Robert Frisch. Great relativist and humanitarian — life and times of Albert Einstein 1879–1955. New Scientist, 81(1145):752– 755, ???? 1979. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Frisch:1979:OLK

[Fri79d] Otto Robert Frisch. Obituary: Lew Kowarski. Nature, 282(5738): 541, November 29, 1979. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v282/n5738/abs/282541a0;   http://www.
nature.com/nature/journal/v282/n5738/pdf/282541a0.pdf.

Frisch:1979:WLR

[Fri79e] Otto Robert Frisch. What Little I Remember. Cambridge University Press, Cambridge, UK, 1979. ISBN 0-521-22297-4. xi + 227 pp. LCCN QC16.F75 A38. URL http://www.loc.gov/

catdir/description/cam022/78018096.html;   http://www.
loc.gov/catdir/toc/cam026/78018096.html.

Frisch:1979:YCP

[Fri79f] Otto Robert Frisch. You can prove anything with statistics. In Duncan and Weston-Smith [DWS79], pages 171–179. ISBN 0-08-021978-0. LCCN AZ999 .L94 1979.

Frisch:1981:WIM

[Fri81] Otto Robert Frisch. Woran ich mich erinnere : Physik und Physiker meiner Zeit. (German) [What I Remember: Physics and Physicists of My Time]. Wissenschaftliche Verlagsgesellschaft, Stuttgart, West Germany, 1981. ISBN 3-8047-0614-2. 276 pp.

Frisch:2009:TPL

[Fri09a] O. R. Frisch. Take a photon . . .. Contemporary Physics, 50(1):59– 67, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:2009:CPQ

[Fri09b] Otto Robert Frisch. The conceptual problem of quantum theory from the experimentalist’s point of view. In Bastin [Bas09], chapter 2, pages 13–21. ISBN 0-521-11548-5, 0-521-07956-X. LCCN QC174.1 .Q83 2009.

Frisch:1931:BMG

[FS31] Otto Robert Frisch and Otto Stern. Beugung von Materiestrahlen. (German) [Diffraction of matter rays]. In Handbuch der Physik, volume 22(2), pages 313–354. ????, ????, 1931.

Frisch:1932:SRM

[FS32] R. Frisch and O. Stern. Die spiegelnde Reflexion von Molekularstrahlen. (German) [The specular reflection of molecular beams]. Die Naturwissenschaften, 20(39):721, September 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Frisch:1933:ERIb

[FS33a] R. Frisch and E. Segr. Über die Einstrahlung der Richtungsquantelung. II. (German) [On the irradiation of directional quantization. II]. Zeitschrift für Physik, 81(5–6):424, May 1933. CODEN ZEPYAA. ISSN 0044-3328.

Frisch:1933:RSQ

[FS33b] R. Frisch and E. Segrè. Ricerche sulla quantizzazione spaziale. (Italian) [Spatial quantization researches]. Il Nuovo Cimento (8), 10(2):78–91, February 1933. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL https://link.springer.

com/article/10.1007/BF02957540.

Frisch:1933:ERIa

[FS33c] R. Frisch and E. Segrè. Über die Einstellung der Richtungsquantelung. II. (German) [On setting the direction quantization. II]. Zeitschrift für Physik, 80(9–10):610–616, September 1933. CO-DEN ZEPYAA. ISSN 0044-3328.

Frisch:1933:ASR

[FS33d] R. Frisch and O. Stern. Anomalien bei der spiegelnden Reflexion und Beugung von Molekularstrahlen an Kristallspaltflächen. I. (German) [Anomalies in specular reflection and diffraction of molecular beams at crystal slits. I]. Zeitschrift für Physik, 84(7– 8):430–442, July 1933. CODEN ZEPYAA. ISSN 0044-3328. URL

https://link.springer.com/article/10.1007/BF01342223.

Frisch:1933:MAW

[FS33e] R. Frisch and O. Stern. Über die magnetische Ablenkung von Wasserstoffmolekülen und das magnetische Moment des Protons. I. (German) [On the magnetic deflection of hydrogen molecules and the magnetic moment of the proton. I]. Zeitschrift für Physik, 85(1–2):4–16, January 1933. CODEN ZEPYAA. ISSN 0044-

3328. URL https://link.springer.com/article/10.1007/
BF01330773.

Frisch:1935:VSN

[FS35] O. R. Frisch and E. T. Sørensen. Velocity of slow neutrons. Nature, 136(3433):258, August 17, 1935. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v136/n3433/pdf/136258a0.pdf.

Fermi:1947:DNM

[FTW47] Enrico Fermi, Edward Teller, and Victor F. Weisskopf. The decay of negative mesotrons in matter. Physical Review, 71(5):314–315, March 1, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://inspirehep.net/record/45186;
 http://prola.aps.org/abstract/PR/v71/i5/p314_1.

Frisch:1937:CSN

[FvK37a] O. R. Frisch, H. von Halban, Jr., and Jørgen Koch. Capture of slow neutrons in light elements. Nature, 140(3551), November 20, 1937. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v140/
n3551/pdf/140895b0.pdf.

Frisch:1937:MFA

[FvK37b] O. R. Frisch, H. von Halban, Jr., and Jørgen Koch. The magnetic field acting upon neutrons inside magnetized iron. Nature, 140 (3539):360, August 28, 1937. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v140/n3539/pdf/140360a0.pdf.

Frisch:1937:MMM

[FvK37c] O. R. Frisch, H. von Halban, Jr., and Jørgen Koch. A method of measuring the magnetic moment of free neutrons. Nature, 139 (3522):756–757, May 1, 1937. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v139/n3522/pdf/139756a0.pdf.

Frisch:1937:SMM

[FvK37d] O. R. Frisch, H. von Halban, Jr., and Jørgen Koch. Sign of the magnetic moment of free neutrons. Nature, 139(3528):1021, June 12, 1937. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v139/n3528/pdf/1391021a0.pdf.

Frisch:1937:SCN

[FvK37e] O. R. Frisch, H. von Halban, Jr., and Jørgen Koch. The slowingdown and capture of neutrons in hydrogenous substances. Math.- fys. Meddr, 15(10):??, ???? 1937.

Frisch:1937:TEC

[FvK37f] O. R. Frisch, H. von Halban, Jr., and Jørgen Koch. Temperature equilibrium of C-neutrons. Nature, 139(3526):922–923, May 29, 1937. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v139/
n3526/pdf/139922b0.pdf.

Frisch:1938:SEM

[FvK38] O. R. Frisch, H. von Halban, Jr., and Jørgen Koch. Some experiments on the magnetic properties of free neutrons. Physical Review, 53(9):719–726, May 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL https://journals.

aps.org/pr/abstract/10.1103/PhysRev.53.719.

Frisch:1967:DFH

[FW67] Otto Robert Frisch and John A. Wheeler. The discovery of fission: How it all began and mechanism of fission. Physics Today, 20 (11):43–52, November 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v20/i11/p43_s1. Reprinted in [WP85,

pages 272–281].

Frisch:1968:DF

[FW68] O. R. Frisch and J. A. Wheeler. The discovery of fission. Physics Bulletin, 19(1):13–20, January 1968. CODEN PHSBB4. ISSN 0031-9112. This article is being published by arrangement with Physics Today. The article first appeared in the November 1967 issue of Physics Today.

Frisch:1985:DF

[FW85] Otto R. Frisch and John A. Wheeler. The discovery of fission. In Weart and Phillips [WP85], pages 272–281. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [FW67].

Graetzer:1971:DNF

[GA71] Hans G. Graetzer and David L. Anderson. The discovery of nuclear fission: a documentary history, volume 20 of Van Nostrand Reinhold momentum books. Van Nostrand Reinhold, New York, NY, USA, 1971. viii + 120 pp. LCCN QC790 .G68.

Gowing:1974:IDBa

[GA74a] Margaret Gowing and Lorna Arnold. Independence and deterrence: Britain and atomic energy, 1945–1952. Macmillan, London, UK, 1974. ISBN 0-333-15781-8 (vol. 1). ???? pp. LCCN QC773.3.G7 G68 1974b.

Gowing:1974:IDBb

[GA74b] Margaret Gowing and Lorna Arnold. Independence and deterrence: Britain and atomic energy, 1945–1952. St. Martin’s Press, New York, NY, USA, 1974. ???? pp. LCCN QC773.3.G7 G68 1974.

Garratt:1963:PSS

[Gar63a] Arthur Garratt, editor. Penguin science survey, 1963-A: astronomy, chemistry, cosmology, education, mathematics, physics, space research. Penguin, London, UK and New York, NY, USA, 1963. 223 pp.

Garrett:1963:DFP

[Gar63b] Alfred Benjamin Garrett. The discovery of the fission process: Lise Meitner and O. R. Frisch. In The flash of genius [Gar63c], pages 227–232. LCCN Q125 .G38.

Garrett:1963:FG

[Gar63c] Alfred Benjamin Garrett. The flash of genius. Van Nostrand, Princeton, NJ, USA, 1963. x + 249 pp. LCCN Q125 .G38.

Gamow:1947:RTM

[GHJ47] George Gamow, Paul R. Heyl, and James J. Jelinek. Review: They move in the void:: Meet the Atoms by O. R. Frisch, The Atomic Story by John W. Campbell, Explaining the Atom by Selig Hecht. The Scientific Monthly, 65(2):165–167, August 1947. CO-DEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic).

URL http://www.jstor.org/stable/pdfplus/19016.pdf.

Ginoux:2023:FPI

[GJ23] Jean-Marc Ginoux and Franck Jovanovic. Frisch’s propagation– impulse model: a comprehensive mathematical analysis. Foundations of Science, 28(1):57–84, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://

link.springer.com/article/10.1007/s10699-021-09827-9.

Goldschmidt:1987:PAF

[Gol87] Bertrand Goldschmidt. Pionniers de l’atome. (French) [Pioneers of the atom]. Stock, Paris, France, 1987. ISBN 2-234-02070-0. 484 + 8 pp. LCCN QC773 .G64 1987. URL http://gallica.

bnf.fr/ark:/12148/bpt6k4803658x. See also English translation

[Gol90].

Goldschmidt:1990:AR

[Gol90] Bertrand Goldschmidt. Atomic Rivals. Rutgers University Press, New Brunswick, NJ, USA, 1990. ISBN 0-8135-1518-1 (hardcover), 0-8135-1519-X (paperback). xvii + 372 + 8 pp. LCCN QC773 .G6413 1990. English translation of [Gol87].

Gowing:1964:BAEb

[Gow64] Margaret Gowing. Britain and Atomic Energy, 1939–1945. St. Martin’s Press, New York, NY, USA, 1964. xvi + 464 pp. LCCN QC773.A1 G6 1964. With an introductory chapter by Kenneth Jay.

Graetzer:1964:DNF

[Gra64] Hans G. Graetzer. Discovery of nuclear fission. American Journal of Physics, 32(9):9–15, January 1964. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). See [HS39, Bey49].

Gregory:2000:BB

[Gre00] David Gregory. Brum and the bomb. BBC News Web story.,

February 17, 2000. URL http://news.bbc.co.uk/2/hi/uk_
news/643913.stm.

Glazer:2015:CCA

[GT15] A. M. (Anthony Michael) Glazer and Patience Thomson, editors. Crystal clear: the autobiographies of Sir Lawrence and Lady Bragg. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-874430-7 (hardcover), 0-19-106179-4 (ebook). xx + 427 pp. LCCN QC16.B66 C79 2015. URL http:

//www.loc.gov/catdir/enhancements/fy1604/2015430192-b.
html;   http://www.loc.gov/catdir/enhancements/fy1604/
2015430192-d.html;   http://www.loc.gov/catdir/enhancements/
fy1604/2015430192-t.html.

Habashi:2009:INM

[Hab09] Fathi Habashi. Ida Noddack and the missing elements. Education in Chemistry, 46(2):48–51, March 2009. CODEN EDCHAU. ISSN 0013-1350 (print), 1749-5326 (electronic).

URL http://www.rsc.org/education/eic/issues/2009March/
ida-noddack-rhenium-nuclear-fission.asp.

Hahn:1958:DF

[Hah58] Otto Hahn. The discovery of fission. Scientific American, 198(??):76–84, February 1958. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

crownedanarchist.com/emc2/discovery_of_fission.doc;
http://www.nature.com/scientificamerican/journal/v198/
n2/pdf/scientificamerican0258-76.pdf.

Hall:1961:BRA

[Hal61] A. Rupert Hall. Book review: Atomic Physics Today, by Otto R. Frisch. Scientific American, 205(6):185, December 1961. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.jstor.org/stable/10.2307/
24937176;   http://www.nature.com/scientificamerican/
journal/v205/n6/pdf/scientificamerican1261-177.pdf.

Hendry:1980:BRR

[Hen80] John Hendry. Book reviews: Reminiscence and the contemporary history of science: What Little I Remember, by Otto Frisch. Nuclear Physics in Retrospect. Proceedings of a Symposium on the 1930s, by Roger H. Stuewer. British Journal for the History of Science, 13(3):258–262, November 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://

www.jstor.org/stable/4026200.

Haslett:1960:SS

[HJ60] A. W. Haslett and John St John, editors. Science Survey. Vista Books, London, UK, 1960. ???? pp.

Heinz:2017:EHS

[HJP17] Andreas Martin Heinz, Björn Jonson, and Imre Pázsit. EPS historic sites: Uddmanska House, Kungälv, Sweden. Europhysics News, 48(1):4–5, ???? 2017. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (electronic). URL https://www.

europhysicsnews.org/articles/epn/pdf/2017/01/epn2017-
48-1.pdf.

Hoffmann:2001:OH

[Hof01] Klaus Hoffmann. Otto Hahn. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-95057-5, 1-4613-0101-7.

Hahn:1939:NVB

[HS39] Otto Hahn and Fritz Strassmann. Über den Nachweis und das Verhalten der bei der Bestrahlung des Urans mittels Neutronen entstehenden Erdalkalimetalle. (German) [Concerning the existence of alkaline earth metals resulting from the neutron irradiation of uranium]. Die Naturwissenschaften, 27(1):11–15, January 1939. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). A facsimile is also available in [Bey49, pages 87–91] and in [Gra64]. Abridged English translation in [GA71, pages 44–47].

Kaempffert:1939:WSW

[Kae39] Waldemar Kaempffert. This week in science: When uranium splits: Doubtful source of power; cancer and x-rays; neutron possibilities. New York Times, ??(??):D9, March 5, 1939. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/102937542.

Kaempffert:1945:SSB

[Kae45] Waldemar Kaempffert. Story of scientists’ ‘battle’ for atom bomb secret revealed in Smyth Report. New York Times, ??(??):8, August 16, 1945. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/107074385/.

Kaempffert:1948:RRB

[Kae48] Waldemar Kaempffert. The revolution that radium began: Fifty years after the Curies’ great discovery, nuclear physics is still a realm unbounded. New York Times, ??(??):SM13, SM25, SM27, December 26, 1948. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/108348269/.

Kockel:1960:BHW

[KED^{+}60] B. Kockel, W. H. Estphal, F. Dessauer, K. Strubecker, H. Siedentopf, F. Borgnis, H. Falkenhagen, W. Finkelnburg, and R. Vieweg. Bücher: Heisenberg: Wandlungen in den Grundlagen der Naturwissenschaft/Heisenberg: Physik und Philosophie/Frisch: Beiträge zur Physik und Chemie des 20. Jahrhunderts/Bellman: Introduction to Matrix Analysis/Mehlin: Astronomy/Born und Wolf: Principles of Optics/Mandl: Introduction to Quantum Field Theory/Marley, Morgan: Health Physics/Churchman: Measurement Definitions and Theories. Physikalische Blätter, 16(6):344– 347, June 1960. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-

3722 (electronic). URL http://onlinelibrary.wiley.com/doi/
10.1002/phbl.19600160610/abstract.

Kelly:2004:RMP

[Kel04] Cynthia C. Kelly, editor. Remembering the Manhattan Project: Perspectives on the Making of the Atomic Bomb and Its Legacy. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2004. ISBN 981-256-040-8. LCCN QC773.A1 R46 2004.

Katzir:2013:TTH

[KLR13] Shaul Katzir, Christoph Lehner, and Jürgen Renn, editors. Traditions and transformations in the history of quantum physics: HQ-3, Third International Conference on the History of Quantum Physics, Berlin, June 28–July 2, 2010, volume 5 of Max Planck research library for the history and development of knowledge. Proceedings. Edition Open Access, Berlin, Germany, 2013. ISBN 3-8442-5134-0. LCCN QC173.98. URL http://www.edition-open-

access.de/proceedings/5/.

Kragh:2014:NPP

[Kra14] Helge Kragh. The names of physics: plasma, fission, photon. European Physical Journal H, 39(3):263–281, September 2014.

CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2014-50007-7.

Kuhn:1963:IOF

[Kuh63] Thomas S. Kuhn. Interview of Otto Frisch. AIP Web site., May

8, 1963. URL https://www.aip.org/history-programs/niels-
bohr-library/oral-histories/4615.

Laurence:1951:DHR

[Lau51] William L. Laurence. Day of Hiroshima recalls atom race: Two leaders in atomic work at Columbia University. New York Times, ??(??):3, August 6, 1951. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/112188459/fulltextPDF.

Lee:2007:REP

[Lee07] Sabine Lee. Rudolf Ernst Peierls. 5 June 1907–19 September 1995. Biographical Memoirs of Fellows of the Royal Society, 53:265–284, 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL http://rsbm.royalsocietypublishing.org/
content/53/265;   http://rsbm.royalsocietypublishing.
org/content/suppl/2009/05/01/53.0.265.DC1.

Leicester:1968:SBC

[Lei68] Henry M. Leicester. Source Book in Chemistry, 1900–1950. Harvard University Press, Cambridge, MA, USA, 1968. ISBN 0-674-36669-7, 0-674-36670-0. xvii + 408 pp. LCCN QD3.

Lightman:2005:DGBb

[Lig05] Alan P. Lightman. The discoveries: great breakthroughs in twentieth-century science. Pantheon, New York, NY, USA, 2005. ISBN 0-676-97789-8, 0-375-42168-8 (hardcover). xviii + 553 + 16 pp. LCCN Q180.55.D57 .L53. URL http://www.loc.gov/

catdir/enhancements/fy0623/2005040854-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0623/2005040854-d.
html;   http://www.loc.gov/catdir/enhancements/fy0623/
2005040854-t.html.

Lindsay:1960:BRT

[Lin60] R. Bruce Lindsay. Book review: Trends in Atomic Physics: essays dedicated to Lise Meitner, Otto Hahn, Max von Laue on the occasion of their 80th birthday. Edited by O. R. Frisch, F. A. Paneth, F. Laves, P. Rosbaud. 285 pp. (Vieweg & Sohn, Germany) Interscience Publishers, Inc., New York, 1959. $7.50. Physics Today, 13(9):45, September 1960. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://scitation.aip.

org/content/aip/magazine/physicstoday/article/13/9/10.
1063/1.3057115.

Lipson:1959:BRB

[Lip59] H. Lipson. Book review: Beiträge Sur Physik Und Chemie Des 20 Jahrhunderts, by O. R. Frisch, F. A. Paneth, F. Laves and P. Rosbaud Braunschweig: Friedr. Vieweg & Sohn. Pp. 285. Physics Bulletin, 10(12):313, 1959. CODEN PHSBB4. ISSN 0031-9112.

URL http://stacks.iop.org/0031-9112/10/i=12/a=016.

Marsden:1954:RML

[Mar54] E. Marsden. The Rutherford Memorial Lecture, 1954. Rutherford — his life and work, 1871–1937. Proceedings of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 226(1166): 283–305, 1954. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-

9169 (electronic). URL http://rspa.royalsocietypublishing.

org/content/226/1166/283. Lecture delivered in South Africa in May 1954.

McCarty:2001:NLH

[McC01] Marilu Hurt McCarty. The Nobel laureates: how the world’s greatest economic minds shaped modern thought. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-135614-2. xvi + 397 pp. LCCN HB87 .M337 2001.

McCauley:2024:HRT

[McC24] Joseph L. McCauley. Heisenberg’s 1939 reactor theory, Serber’s 1943 Los Alamos Primer, and Heisenberg’s 1945 Farm Hall critical mass calculation. American Journal of Physics, 92(10):765–774, October 2024. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). See comments and corrections [Ree24a].

McCauley:2025:PCR

[McC25] Joseph L. McCauley. Predictions of critical radii for reactors and bombs 1939–45 including the Frisch–Peierls memorandum. European Physical Journal H, 50(1):??, December 2025. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-024-00088-1.

Meitner:1962:RWR

[Mei62] Lise Meitner. Right and wrong roads to the discovery of nuclear energy. International Atomic Energy Agency Bulletin, 4(0):4–6, ???? 1962. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-

2690 (electronic). URL http://www.iaea.org/Publications/
Magazines/Bulletin/Bull040su/04004790608su.pdf.

Meitner:1939:DUN

[MF39a] Lise Meitner and Otto Robert Frisch. Disintegration of uranium by neutrons: a new type of nuclear reaction. Nature, 143(3615): 239–240, February 11, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v143/n3615/pdf/143239a0.pdf. This pa-

per, and [MF39c], both submitted 16 January 1939 (see [Mei62]), provided the first published explanation of nuclear disintegration, called ‘nuclear fission’ by Frisch, that was first observed experimentally by Hahn and Strassmann in December 1939 [HS39]. Frisch’s paper [Fri39c] describes the first experimental confirmation. It was these results that Niels Bohr intended to hold confidential until their journal publication during his January 1939 trip to the USA, but his traveling companion Léon Rosenfeld [Ros72] wasn’t informed of that intent, and the news escaped and spread quickly.

Meitner:1939:PFUa

[MF39b] Lise Meitner and Otto Robert Frisch. On the products of the fission of uranium and thorium under neutron bombardment. Math.- fys. Meddr, 17(5):1–13, 1939.

Meitner:1939:PFUb

[MF39c] Lise Meitner and Otto Robert Frisch. Products of the fission of the uranium nucleus. Nature, 143(3620):471–472, March 18, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v143/
n3620/pdf/143471a0.pdf. See note in [MF39a].

Meitner:1968:DUN

[MF68] Lise Meitner and Otto Robert Frisch. Disintegration of uranium by neutrons. A new type of nuclear reaction. In Source Book in Chemistry, 1900–1950 [Lei68], pages 244–246. ISBN 0-674-36669-7, 0-674-36670-0. LCCN QD3. URL https:

//www.degruyter.com/view/books/harvard.9780674366701/
harvard.9780674366701.c80/harvard.9780674366701.c80.xml.

Meitner:1969:DUN

[MF69] Lise Meitner and O. R. Frisch. Disintegration of uranium by neutrons: a new type of nuclear reaction (reprinted from Nature, February 11, 1939). Nature, 224(5218):466–467, November 1, 1969. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v224/
n5218/abs/224466a0.html.

Milonni:2009:TPL

[Mil09] Dr. P. W. Milonni. Take a photon … (1965) by O. R. Frisch. Contemporary Physics, 50(1):57, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mitchum:1986:BRB

[Mit86] Carl Mitchum. Book review: A Theory of Technology: Continuity and Change in Human Development by Thomas R. DeGregori; International Technology Transfer: Concepts, Measures, and Comparisons by Nathan Rosenberg; Claudio Frischtak. Isis, 77 (3):552–554, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231653.

Mladjenovic:1998:DYN

[Mla98] Milorad Mladjenović. The Defining Years in Nuclear Physics, 1932–1960s. IOP Publishing, Bristol, UK, 1998. ISBN 0-7503-0472-3 (hardcover). xx + 441 pp. LCCN QC773 .M54 1998.

Medawar:1999:HGS

[MP99] Jean S. Medawar and David Pyke. Hitler’s Gift: Scientists Who Fled Nazi Germany. Richard Cohen Books and the European Jewish Publication Society, London, UK, 1999. ISBN 1-86066-172-6. xx + 268 pp. LCCN Q141 .M385 1999.

Medawar:2001:HGT

[MP01] J. S. Medawar and David Pyke. Hitler’s Gift: the True Story of the Scientists Expelled by the Nazi Regime. Arcade Publishing, New York, NY, USA, 2001. ISBN 1-55970-564-7. xx + 268 pp. LCCN Q141 .M385 2001. Foreword by Max Perutz.

Muller:1968:DF

[Mül68] Walter Müller. The discovery of fission. Physics Bulletin, 19(5): 157, May 1968. CODEN PHSBB4. ISSN 0031-9112.

Neffe:2005:EBG

[Nef05] Jürgen Neffe. Einstein: eine Biographie. (German) [Einstein: a Biography]. Rowohlt, Reinbek bei Hamburg, Germany, second edition, 2005. ISBN 3-498-04685-3. 490 + 14 pp. LCCN QC16.E5 N44 2005.

Neffe:2006:EBG

[Nef06] Jürgen Neffe. Einstein: eine Biographie. (German) [Einstein: a Biography], volume 61937 of rororo. Rowohlt Taschenbuch Verlag, Reinbek bei Hamburg, Germany, 2006. ISBN 3-499-61937-7 (paperback). 490 + 8 pp. LCCN ???? EUR 9.90, SFR 17.90.

Neffe:2007:EB

[Nef07] Jürgen Neffe. Einstein: a biography. Farrar, Strauss, and Giroux, New York, NY, USA, 2007. ISBN 0-374-14664-0 (hardcover). x + 461 + 16 pp. LCCN QC16.E5 N4413 2007. URL http://www.loc.

gov/catdir/enhancements/fy0705/2006026136-b.. Translated

by Shelley Frisch from the original German edition Einstein: eine Biographie (ISBN 3-498-04685-3, 3-499-61937-7) [Nef05, Nef06].

Noddack:1934:EGE

[Nod34] Ida Noddack. Über das Element 93. (German) [On element 93 [neptunium]]. Zeitschrift für Angewandte Chemie, 47(37):653– 655, September 15, 1934. ISSN 0932-2132. URL http://en.

wikipedia.org/wiki/Neptunium;   http://onlinelibrary.
wiley.com/doi/10.1002/ange.19340473707/abstract;   http:
//www.chemteam.info/Chem-History/Noddack-1934.html. Ac-

cording to Frisch [FW67, page 48], “Ida Noddack, a German chemist, quite rightly pointed out that they might be lighter elements [after bombardment of uranium by neutrons]; but her comments (published in a journal not much read by chemists and hardly at all by physicists) were regarded as mere pedantry. She did not indicate how such light elements could be formed; her paper had probably no effect whatever on later work.”. English translation in [GA71, pages 16–20].

Oliphant:1966:TEa

[Oli66a] Mark L. Oliphant. The two Ernests — I. Physics Today, 19 (9):35–49, September 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://scitation.aip.

org/content/aip/magazine/physicstoday/article/19/9/10.

1063/1.3048466. Reprinted in [Oli85].

Oliphant:1966:TEb

[Oli66b] Mark L. Oliphant. The two Ernests — II. Physics Today, 19(10):41–51, October 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://scitation.

aip.org/content/aip/magazine/physicstoday/article/19/
10/10.1063/1.3047765. Reprinted in [Oli85].

Oliphant:1985:TE

[Oli85] Mark L. Oliphant. The two Ernests. In Weart and Phillips [WP85], pages 173–193. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [Oli66a, Oli66b].

Olsen:1963:TGS

[Ols63] Arthur J. Olsen. Trackdown of the German scientist: Nazism and defeat scattered the leaders of a once-great scientific establishment. Herewith a review of where they went and what some of them did when they got there. New York Times, ??(??): 214, September 22, 1963. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/116463805/.

Oliphant:1938:RSA

[OPM38] M. L. Oliphant, R. Peierls, and P. B. Moon. Radioactivity and subatomic phenomena. Annual Reports on the Progress of Chemistry, 35:7–35, 1938. CODEN ARPCAW. ISSN 0365-6217 (print), 1754-7512 (electronic).

Pearson:2024:CFP

[Pea24] J. M. Pearson. Comments on the Frisch–Peierls estimate of the critical mass of a uranium fission bomb. Nuclear Technology, ?? (??):1–5, February 2024. CODEN NUTYBB. ISSN 1943-7471 (print), 1943-7471 (electronic).

Peierls:1979:BRP

[Pei79] Rudolf Peierls. Book review: A physicist who enjoys it. What Little I Remember, by Otto R. Frisch, pp. 219. (Cambridge University Press: Cambridge, 1979). Nature, 280(5719):257–259, July 19, 1979. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1979Natur.280..257P.

Peierls:1980:ORF

[Pei80] Rudolf Peierls. Obituary: O. R. Frisch, 1904–1979. Nature, 284 (5752):196–197, March 13, 1980. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1980Natur.284..196P.

Peierls:1981:ORF

[Pei81] Sir Rudolf Peierls, F.R.S. Otto Robert Frisch, 1 October 1904–22 September 1979. Biographical Memoirs of Fellows of the Royal Society, 27:283–306, November 1981. CODEN BM-FRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL

http://www.jstor.org/stable/10.2307/769874;   http://
www.jstor.org/stable/769874;   https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1981.0012.

Peierls:1986:FFW

[Pei86] Rudolf Peierls. Fact or fission [wild speculation]. London Review of Books, 8(18):4, October 23, 1986. ISSN 0260-9592. URL https:

//www.lrb.co.uk/v08/n18/letters#letter4.

Peierls:1991:MST

[Pei91] Rudolf Peierls. More Surprises in Theoretical Physics. Princeton Series in Physics. Princeton University Press, Princeton, NJ, USA, 1991. ISBN 0-691-08576-5 (hardcover), 0-691-02522-3 (paperback). viii + 106 pp. LCCN QC20 .P345 1991. URL http:/

/www.loc.gov/catdir/description/prin021/90023189.html;
 https://ui.adsabs.harvard.edu/#abs/1991mstp.book.....
P.

Peierls:1997:AH

[Pei97] Sir Rudolf Ernst Peierls. Atomic Histories, volume 18 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1997. ISBN 1-56396-243-8 (hardcover). xvii + 378 pp. LCCN QC71 .P38 1997.

Perrin:1939:CRC

[Per39] Francis Perrin. Calcul relatif aux conditions éventuelles de transmutation en chaîne de l’uranium. (French) [Calculation of possible conditions on the uranium chain transmutation]. Comptes rendus de l’Académie des sciences, Paris, 208(18):1394–

1396, May 1, 1939. URL http://gallica.bnf.fr/ark:/12148/
bpt6k6238836c/f26.image.r=perrin%20calcul%20relative%
20aux%20conditions#.

Peierls:1940:EFNb

[PF40] Rudolf Peierls and Otto R. Frisch. The effects of fast neutrons in ordinary uranium. Report MS-4A (PRO AB4/832), Directorate of Tube Alloys, ????, UK, 1940.

Pohl:1962:ED

[Poh62] Frederik Pohl. The expert dreamers. Doubleday, New York, NY, USA, 1962. 248 pp. LCCN PZ1 .P745 Ex. Reprinted in 1968 by Avon Books, New York.

Peierls:2004:FOR

[PR04] Rudolf Peierls and Xavier Roqué. Frisch, Otto Robert (1904– 1979), physicist. In Oxford Dictionary of National Biography, page ?? Oxford University Press, Walton Street, Oxford OX2

6DP, UK, 2004. URL https://www.oxforddnb.com/search?q=
Frisch%2C+Otto+Robert+%281904%E2%80%931979%29%2C+physicist.

Prasad:1993:RRR

[Pra93] Rajendra Prasad. Review: Re-reading classics A Mathematician’s Apology by G. H. Hardy; What Little I Remember by Otto Frisch; What Is Life?, with Mind and Matter and Autobiographical Sketches by Erwin Schrödinger; Mr. Tompkins in Paperback by George Gamow. Social Scientist, 21(3/4):98–101, April 1993. CODEN ???? ISSN 0970-

0293. URL http://www.jstor.org/stable/10.2307/3517634;
 http://www.jstor.org/stable/pdfplus/3517634.pdf.

Reed:2015:ABS

[Ree15a] Bruce Cameron Reed. The atomic bomb: the story of the Manhattan Project: how nuclear physics became a global geopolitical game-changer. IOP concise physics. Morgan and Claypool Publishers and IOP Publishing, San Rafael, CA, USA and Bristol, UK, 2015. ISBN 1-62705-990-3 (print), 1-62705-991-1 (e-book), 1-62705-993-8 (mobi). ISSN 2053-2571 (print), 2054-7307 (electronic). 239 (est.) pp. LCCN QC773.3.U5 R443 2015eb. URL

http://iopscience.iop.org/book/978-1-6270-5991-6.

Reed:2015:BS

[Ree15b] Bruce Cameron Reed. The background science. In The atomic bomb: the story of the Manhattan Project: how nuclear physics became a global geopolitical game-changer [Ree15a], pages 2:1–2:30.

ISBN 1-62705-990-3 (print), 1-62705-991-1 (e-book), 1-62705-993-8 (mobi). ISSN 2053-2571 (print), 2054-7307 (electronic). LCCN

QC773.3.U5 R443 2015eb. URL http://iopscience.iop.org/
book/978-1-6270-5991-6.

Reed:2022:CPF

[Ree22] B. Cameron Reed. Comments on the physics of the Frisch–Peierls memorandum. Nuclear Technology, 208(12):1890–1893, August 2022. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Reed:2024:HB

[Ree24a] B. Cameron Reed. On Heisenberg and the bomb. American Journal of Physics, 92(12):908–909, December 2024. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). This short paper corrects some errors in [McC24] and supplies additional historical material.

Reed:2024:RFP

[Ree24b] B. Cameron Reed. Revisiting the Frisch–Peierls memorandum. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-024-00070-x.

Reines:1972:CFOa

[Rei72] Frederick Reines, editor. Cosmology, fusion and other matters: George Gamow memorial volume. Colorado Associated University Press, Boulder, CO, USA, 1972. ISBN 0-87081-025-1. xiv + 320 pp. LCCN QC780 .C65.

Reinhardt:2007:CSC

[Rei07] Carsten Reinhardt. Chemical sciences in the 20th century: bridging boundaries. Wiley-VCH, New York, NY, USA, 2007. ISBN 1-281-84243-5, 3-527-30271-9, 3-527-61273-4 (e-book). xviii + 281 pp. LCCN QD15 .C44 2001. URL http://sfx.ethz.

ch/sfx\_locater?sid=ALEPH:EBI01\%26genre=book\%26isbn=
9783527302710\%26id=doi:10.1002/9783527612734.

Rife:2006:LMD

[Rif06] Patricia Rife. Lise Meitner and the dawn of the nuclear age. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2006. ISBN 0-8176-4559-4 (softcover), 0-8176-3732-X (hardcover). ???? pp. LCCN ???? URL http:

//www.loc.gov/catdir/enhancements/fy0812/2006935949-d.
html;   http://www.loc.gov/catdir/enhancements/fy0812/

2006935949-t.html. Foreword by John Archibald Wheeler.

Rosenfeld:1972:NR

[Ros72] Léon Rosenfeld. Nuclear reminiscences. In Reines [Rei72], pages 289–299. ISBN 0-87081-025-1. LCCN QC780 .C65. URL http:/

/adsabs.harvard.edu/abs/1972cht..conf..289R.

Rose:1998:HNA

[Ros98] Paul Lawrence Rose. Heisenberg and the Nazi atomic bomb project: a study in German culture. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1998. ISBN 0-520-21077-8, 0-585-32190-6 (e-book). xx + 352 pp. LCCN

QC16.H35 R67 1998. URL http://preterhuman.net/texts/
religion.occult.new_age/occult.conspiracy.and.related/
Rose,%20Paul%20Lawrence%20-%20Heisenberg%20and%20the%
20Nazi%20Atomic%20Bomb%20Project.pdf;   http://www.scribd.
com/doc/176628605/Rose-Paul-Lawrence-Heisenberg-and-the-
Nazi-Atomic-Bomb-Project.

Rotblat:1954:AES

[Rot54] Joseph Rotblat, editor. Atomic energy: a survey; [being news and views of atomic energy given as a course of lectures at the University of London … during January and February 1954]. Taylor & Francis, London, UK, 1954. viii + 71 pp. Published for Atomic Scientist’ Association.

Rotblat:2000:PJR

[Rot00] Joseph Rotblat. My early years as a physicist in Poland: a talk given to the Group on Monday 8th March 1999. IOP History of Physics Group Newsletter, 13:9–23, Spring 2000. ISSN 1756-

168X. URL http://www.iop.org/activity/groups/subject/
hp/newsletter/archive/file_66506.pdf.

Rowlands:2009:LC

[Row09] Peter Rowlands. The Liverpool cyclotrons. IOP History of Physics Newsletter, 25:31–46, February 2009. ISSN 1756-

168X. URL http://www.iop.org/activity/groups/subject/
hp/newsletter/archive/file_64509.pdf.

Rozental:1967:NBH

[Roz67] S. (Stefan) Rozental, editor. Niels Bohr: his life and work as seen by his friends and colleagues. North-Holland, Amsterdam, The Netherlands, 1967. 355 pp. LCCN QC16.B63 N53.

Rosenfeld:1964:NBH

[RRK^{+}64] Léon Rosenfeld, Erik Rüdinger, Oskar Klein, Werner Heisenberg Hendrik G. B. Casimir, Otto Robert Frisch, Stefan Rozental, Aage Bohr, Abraham Pais, Jørgen Kalckar, Christian Møller, Mogens Pihl, Viktor F. Weisskopf, Johannes Pedersen, Viggo Kampmann, Richard Courant, Paul A. M. Dirac, Hans Henrik Koch, William Scharff, Mogens Andersen, Hans Bohr, and Niels Bohr. Niels Bohr: Hans liv or virke fortalt af en kreds af venner og medarbejdere. (Danish) [Niels Bohr: His life and works told by a group of friends and co-workers]. J. H. Schultz Forlag, Copenhagen, DK, 1964. 341 pp. LCCN ????

Rapport:1964:P

[RW64] Samuel Rapport and Helen Wright, editors. Physics. New York University Press, New York City, NY, USA, 1964. xiii + 333 pp.

Schmidt-Bocking:2016:OSVa

[SBRT^{+}16a] Horst Schmidt-Böcking, Karin Reich, Alan Templeton, Wolfgang Trageser, and Volkmar Vill, editors. Otto Sterns Veröffentlichungen — Band 1, volume 1. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-662-46953-7.

Schmidt-Bocking:2016:OSVb

[SBRT^{+}16b] Horst Schmidt-Böcking, Karin Reich, Alan Templeton, Wolfgang Trageser, and Volkmar Vill, editors. Otto Sterns Veröffentlichungen — Band 2, volume 2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-662-46962-6.

Schmidt-Bocking:2016:OSVc

[SBRT^{+}16c] Horst Schmidt-Böcking, Karin Reich, Alan Templeton, Wolfgang Trageser, and Volkmar Vill, editors. Otto Sterns Veröffentlichungen — Band 3, volume 3. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-662-46960-X.

Schmidt-Bocking:2016:OSVd

[SBRT^{+}16d] Horst Schmidt-Böcking, Karin Reich, Alan Templeton, Wolfgang Trageser, and Volkmar Vill, editors. Otto Sterns Veröffentlichungen — Band 4, volume 4. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-662-46964-2.

Schwarzschild:2000:BHS

[Sch00] Bertram Schwarzschild. Bohr–Heisenberg Symposium marks Broadway opening of Copenhagen. Physics Today, 53(5):51–52, May 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v53/i5/
p51/s1.

Segre:1985:HPR

[Seg85] Emilio Segrè. Historical perspective: Refugee scientists and nuclear energy. Annals of the New York Academy of Sciences, 452(1): xv–xix, 1985. CODEN ANYAA9. ISBN 0-89766-298-9, 0-89766-299-7 (paperback). ISSN 0077-8923 (print), 1749-6632 (electronic). Sixth International Conference on Collective Phenomena: reports from the Moscow Refusnik Seminar / edited by Inga Fischer-Hjalmars and Joel L. Lebowitz. Contributions from the Moscow Refusnik Seminar and from two International Conferences on Collective Phenomena, one held in Stockholm, Sweden, 1–2 December 1983, and the other in Tel Aviv, Israel, 31 May–1 June 1984.

Segre:1993:MAM

[Seg93] Emilio Segrè. A mind always in motion: the autobiography of Emilio Segrè. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1993. ISBN 0-520-07627-3. xii + 332 pp. LCCN QC16.S35 A3 1993. URL http://ark.

cdlib.org/ark:/13030/ft700007rb;   http://www.loc.gov/
catdir/description/ucal041/92010722.html.

Seidel:1980:BRB

[Sei80] Robert W. Seidel. Book review: What Little I Remember by Otto R. Frisch. Isis, 71(3):517–518, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.2307/230165;
 http://www.jstor.org/stable/230165.

Stein:1945:DNU

[SFFdH45] P. R. Stein, Otto Robert Frisch, Bernard Taub Feld, and F. de Hoffman. Delayed neutrons from U^{235} after short irradiation. Report AECD-1971, U.S. Atomic Energy Commission, Oak Ridge, TN, USA, April 6, 1945. 20 pp. URL https://babel.

hathitrust.org/cgi/pt?id=mdp.39015086446955. Declassified

17 May 1948. Title page undated, but date found in footnote on first document page.

Siegel:1980:BRW

[Sie80] Daniel M. Siegel. Book review: What Little I Remember, by O. R. Frisch. Physics Today, 33(4):56, April 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Sime:1996:MFG

[Sim96] Ruth Lewin Sime. Meitner, Frisch got to fission theory first — the rest found it in Nature. Physics Today, 49(7):92, July 1996. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Sime:2007:SAE

[Sim07] Ruth Lewin Sime. The search for artificial elements and the discovery of nuclear fission. In Chemical sciences in the 20th century: bridging boundaries [Rei07], pages 146–159. ISBN 1-281-84243-5, 3-527-30271-9, 3-527-61273-4 (e-book). LCCN QD15

.C44 2001. URL https://onlinelibrary.wiley.com/doi/pdf/
10.1002/9783527612734.ch08.

Sime:2012:PFO

[Sim12] Ruth Lewin Sime. The politics of forgetting: Otto Hahn and the German Nuclear-Fission Project in World War II. Physics in Perspective (PIP), 14(1):59–94, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-011-0065-6;
http://www.springerlink.com/content/k12202vg92147h68/.

Sime:2014:SPD

[Sim14] Ruth Lewin Sime. Science and politics: The discovery of nuclear fission 75 years ago. Annalen der Physik (1900), 526(3–4): A27–A31, April 2014. ISSN 0003-3804 (print), 1521-3889 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.
1002/andp.201400805.

Smith:1961:RMP

[Smi61] Malcolm K. Smith. Review: Modern physics for laymen: The Atom and Its Nucleus by George Gamow; Atomic Physics Today by Otto R. Frisch. Science, 134(3474):276, July 28, 1961. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://www.jstor.org/stable/pdfplus/1708183.pdf.

Smolin:2007:OEB

[Smo07] Lee Smolin. The other Einstein: Einstein: His Life and Universe by Walter Isaacson. Einstein: A Biography by Jürgen Neffe, translated from the German by Shelley Frisch. Subtle Is the Lord: The Science and the Life of Albert Einstein by Abraham Pais. The Private Lives of Albert Einstein by Roger Highfield and Paul Carter. Einstein in Love: A Scientific Romance by Dennis Overbye. Einstein’s Clocks, Poincaré’s Maps: Empires of Time by Peter Galison. Einstein on Politics edited by David Rowe and Robert Schulmann. Einstein on Race and Racism by Fred Jerome and Rodger Taylor. The Collected Papers of Albert Einstein by Albert Einstein. New York Review of Books, 54(10):76– ??, June 14, 2007. ISSN 0028-7504 (print), 1944-7744 (electronic).

URL http://www.nybooks.com/articles/archives/2007/jun/
14/the-other-einstein/.

Serber:1992:APF

[SR92] R. (Robert) Serber and Richard Rhodes. The Los Alamos primer: the first lectures on how to build an atomic bomb. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1992. ISBN 0-520-07576-5. xxxiii + 98 + 8 pp. LCCN QC773.A1 S47 1992. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http://
www.loc.gov/catdir/bios/ucal051/91014068.html;   http:
//www.loc.gov/catdir/description/ucal041/91014068.html.

Stanley:2008:BRE

[Sta08] Matthew Stanley. Book reviews: Einstein: Essence or Explanation? Einstein: His Life and Universe, by Walter Isaacson. New York: Simon & Schuster, 2007. xxii + 675 pp., illus., index. $32 (cloth). Einstein: A Biography, by Jürgen Neffe. Trans. Shelley Frisch. New York: Farrar, Straus, and Giroux, 2007. x + 461 pp., illus., index. $30 (cloth). Einstein on Politics. David E. Rowe and Robert Schulmann, ed. Princeton: Princeton University Press, 2007. xxxiv + 523 pp., illus., index. $29.95 (cloth). Historical Studies in the Natural Sciences, 38(1):153–161, February 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Stolz:1989:OHL

[Sto89] Werner Stolz. Otto Hahn/Lise Meitner. Vieweg + Teubner Verlag, ????, 1989.

Stuewer:1979:NPR

[Stu79] Roger H. Stuewer, editor. Nuclear physics in retrospect: proceedings of a symposium on the 1930s. University of Minnesota Press, Minneapolis, MN, USA, 1979. ISBN 0-8166-0869-5. LCCN QC773 .S95 1977.

Stuewer:1985:BNF

[Stu85] Roger H. Stuewer. Bringing the news of fission to America. Physics Today, 38(10):48–56, October 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

www.physicstoday.org/resource/1/phtoad/v38/i10/p48_s1.

Stuewer:1994:OLD

[Stu94] Roger H. Stuewer. The origin of the liquid-drop model and the interpretation of nuclear fission. Perspectives on Science, 2 (1):76–129, Spring 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://alsos.wlu.edu/

information.aspx?id=1736. See errata [Ano94].

Stuewer:2013:ACM

[Stu13] Roger H. Stuewer. An act of creation: the Meitner–Frisch interpretation of nuclear fission. In Katzir et al. [KLR13], pages 231–245. ISBN 3-8442-5134-0. LCCN QC173.98. URL http://

www.edition-open-access.de/proceedings/5/.

Stuewer:2018:AIN

[Stu18] Roger H. Stuewer. The Age of Innocence: Nuclear Physics Between the First and Second World Wars. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-186658-X, 0-19-882787-3 (hardback), 0-19-256290-8 (e-book). xv + 484 pp. LCCN QC773 .S78 2018.

Szasz:1992:BSMa

[Sza92a] Ferenc Morton Szasz. British scientists and the Manhattan Project: the Los Alamos years. Macmillan, Basingstoke, UK, 1992. ISBN 1-349-12731-0, 0-333-56597-5. xx + 167 pp. LCCN

???? URL http://link.springer.com/10.1007/978-1-349-
12731-3.

Szasz:1992:BSMb

[Sza92b] Ferenc Morton Szasz. British scientists and the Manhattan Project: the Los Alamos years. St. Martin’s Press, New York, NY, USA, 1992. ISBN 0-312-06167-6. xx + 167 pp. LCCN QC773.3.U5

S97 1991. URL http://www.loc.gov/catdir/enhancements/
fy0809/91019904-d.html.

Tromel:2009:VEK

[Trö09] Martin Trömel. Die vier Entdecker der Kernspaltung. (German) [The four discovers of nuclear fission]. Nachrichten aus der Chemie, 57(1):57, January 2009. CODEN NACHFB. ISSN 1439-9598 (print), 1868-0054 (electronic). URL https://www.

gdch.de/fileadmin/downloads/Publikationen/Nachrichten_
aus_der_Chemie/PDFs/kernspaltung.pdf. See [Ell08].

Walker:2003:OHV

[Wal03] Mark Walker. Otto Hahn: Verantwortung und Verdrängung. (German) [Otto Hahn: responsibility and repression]. Forschungsprogramm ,,Geschichte der Kaiser-Wilhelm-Gesellschaft im Nationalsozialismus” (Research Program “History of the Kaiser Wilhelm Society in the National Socialist Era”), Berlin, Germany,

2003. 64 pp. LCCN ???? URL www.mpiwg-berlin.mpg.de/KWG/
Ergebnisse/Ergebnisse10.pdf.

Walker:2006:OHR

[Wal06] Mark Walker. Otto Hahn: Responsibility and repression. Physics in Perspective (PIP), 8(2):116–163, May 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-006-0277-3.

Weiner:1979:OHO

[Wei79] Charles Weiner. Oral histories; Otto Frisch. American Institute of Physics (AIP) interview., May 3, 1979. URL https:/

/www.aip.org/history-programs/niels-bohr-library/oral-
histories/4616.

West:2004:MCG

[Wes04] Nigel West. Mortal crimes: the greatest theft in history: Soviet penetration of the Manhattan Project. Enigma Books, New York, NY, USA, 2004. ISBN 1-929631-29-4 (hardcover), 1-936274-82-5 (e-book). xxii + 279 + 16 pp. LCCN UB271.R9 W47 2004.

Wheeler:2009:MF

[Whe09] John A. Wheeler. Mechanism of fission. Physics Today, 62(4): 35–38, April 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Winterberg:1996:MHN

[WHF^{+}96] Friedwardt Winterberg, Günter Herrmann, Igor Fodor, Lincoln Wolfenstein, and Mark E. Singer. More on how Nazi Germany failed to develop the atomic bomb. Physics Today, 49(1):11–83, January 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Wilson:1975:AOT

[Wil75] Jane Wilson, editor. All in our time: the reminiscences of twelve nuclear pioneers. Bulletin of the Atomic Scientists, Chicago, IL, USA, 1975. 236 pp. LCCN QC773.A1 A44. Reprinted by the Educational Foundations for Nuclear Science, Chicago, IL, USA (1975).

Wilson:1980:BRW

[Wil80] R. R. Wilson. Book review: What Little I Remember, Otto R. Frisch. Science, 207(4435):1068, March 7, 1980. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.jstor.org/stable/10.2307/1684108.

Ward:2015:RWF

[WM15] Tim Ward and Domenic Mastrippolito. Race for the world’s first atomic bomb: A thousand days of fear. [Los Alamos National] Laboratory produced documentary on the Man-

hattan Project., 2015. URL http://www.governmentattic.
org/11docs/NNSAfilmVideoList_2014.pdf;   https://www.
youtube.com/watch?v=xwpgmEvlRpM. Written by Michael Wil-

son. The hour-long program features interviews with some of those who worked on the Project Hans Bethe, Norman Ramsey, Phil Morrison, Edward Teller and others. The program profiles the creation of the Manhattan Project the obstacles that were faced, life in the “secret city,’ and events that led up to and followed the Trinity Test.

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). 375 pp. LCCN QC7 .H694 1985.